### v1

# GenCO: Generating Diverse Solutions to Design Problems with Combinatorial Nature

Abstract
Generating diverse objects (e.g., images) using generative models (such as GAN or VAE) has achieved impressive results in the recent years, to help solve many design problems that are traditionally done by humans. Going beyond image generation, we aim to find solutions to more general design problems, in which both the diversity of the design and conformity of constraints are important. Such a setting has applications in computer graphics, animation, industrial design, material science, etc, in which we may want the output of the generator to follow discrete/combinatorial constraints and penalize any deviation, which is non-trivial with existing generative models and optimization solvers. To address this, we propose GenCO, a novel framework that conducts end-to-end training of deep generative models integrated with embedded combinatorial solvers, aiming to uncover high-quality solutions aligned with nonlinear objectives. While structurally akin to conventional generative models, GenCO diverges in its role - it focuses on generating instances of combinatorial optimization problems rather than final objects (e.g., images). This shift allows finer control over the generated outputs, enabling assessments of their feasibility and introducing an additional combinatorial loss component. We demonstrate the effectiveness of our approach on a variety of generative tasks characterized by combinatorial intricacies, including game level generation and map creation for path planning, consistently demonstrating its capability to yield diverse, high-quality solutions that reliably adhere to user-specified combinatorial properties.
1 Introduction
Generating diverse and realistic objects with combinatorial properties is an important task with many applications. For example, in video game level design, we may want to generate a variety of levels that are both realistic and valid/playable (Zhang et al., 2020). Here, “valid” may refer to certain discrete characteristics of the level that must be satisfied (e.g., a minimum number of enemies, a path between the level entrance and exit, etc.). In automatic device design, we want to generate a variety of devices that meet foundry manufacturing constraints and optimize physics-related objectives (Schubert et al., 2022). In the design of new molecules, we want to generate a variety of chemically valid molecules with specific properties (Pereira et al., 2021). In all of these examples, the goal is to generate a variety of combinatorial solutions that are both feasible and high quality with respect to a given nonlinear objective. Another situation to consider is when the generated solutions do not have to strictly follow combinatorial constraints, but we want to discourage certain things as a penalty. For example, if we are creating various images of a map, we might want to ensure that the path planning in those maps is efficient (see section 4.2).
The difficulty in these settings arises from the fact that the generated objects need to satisfy specific discrete properties incorporated either as combinatorial constraints or penalties and be of high quality as evaluated by a nonlinear objective. The combinatorial constraints make it challenging to apply gradient-based methods to train standard generative models, such as GAN (Goodfellow et al., 2014), or VAE (Kingma & Welling, 2013), which may generate examples that violate the user-specified constraints. Additionally, the nonlinear objective makes it difficult to use traditional general-purpose combinatorial optimization tools since mixed-integer nonlinear programming (MINLP) is often slow in practice compared to the more heavily researched mixed integer linear programming (MILP), often requiring problem-specific solvers or leveraging specialized structure. Furthermore, traditional optimization solvers are not geared towards generating diverse solutions and do not have the flexibility of deep learning models to quickly adapt to slightly modified settings. Recent work has attempted to address these issues by first training a generative model and then fixing/postprocessing the output after the fact (Zhang et al., 2020). However, these approaches are limited in that they only use the combinatorial solver to fix the output of the generative model and do not use the combinatorial solver in training. As a result, the generative model is not trained with the combinatorial solver in mind (i.e., end-to-end). Thus, the generator may fail to capture the distribution of high-quality feasible solutions as many of the generated examples may be “fixed” to a single solution. Additionally, some work constrains generative model outputs by penalizing the generator based on constraint violation (Chao et al., 2021). While this approach may work in some settings where the constraints must be satisfied, it is unclear how to extend this approach to more complex constraints, such as logical or general combinatorial constraints that can be readily expressed in MILP.
As a result, we propose GenCO that generates diverse solutions to design problems with combinatorial nature. GenCO combines the flexibility of deep generative models with the combinatorial efficiency of optimization solvers. Our approach mainly deals with two cases: (1) when the resulting solution needs to satisfy certain hard combinatorial constraints (hard constraints), and (2) the imposition of a desired penalty if the solutions violate the constraints (soft constraints). To achieve these two goals, we introduce a general concept known as the “combinatorial loss”, which serves as an additional “restriction” on the generated solutions. In the context of combinatorial constraints, the combinatorial loss takes the form of a projection expressed as a mixed-integer linear program (MILP). In the second scenario, this general concept manifests as an extra penalty term within the objective function. This, coupled with the diversity mechanisms inherent in deep generative models, fine-tunes the generated solutions. This process both encourages diversity and ensures adherence to the specified combinatorial properties.
Our main contributions involve introducing a framework, GenCO, that serves two key purposes: 1) generating a range of solutions to nonlinear problems, ensuring they adhere to combinatorial properties; 2) integrating a combinatorial solver into the deep generative learning process. We showcase the effectiveness of our approach through a series of experiments on various generative combinatorial optimization problems. This includes tasks like game level design (section 4.1) and generating maps for path planning (section 4.2).
2 Related Work
The interaction between generative models and combinatorial optimization has seen increased research interest as practitioners seek to integrate or draw inspiration from both paradigms. A prominent line of research has approached the problem of generating objects with combinatorial optimization in mind.
Traditional constrained object generation
Traditional constraint optimization methods were modified to search for multiple feasible solutions for problems concerning building layout (Bao et al., 2013), structural trusses (Hooshmand & Campbell, 2016), networks (Peng et al., 2016), building interiors (Wu et al., 2018), and urban design (Hua et al., 2019). Additionally, an approach based on Markov chain Monte Carlo (Yeh et al., 2012) samples objects that satisfy certain constraints. These approaches employ traditional sampling and optimization methods such as mathematical programming and problem-specific heuristics to obtain multiple feasible solutions. These methods often obtain multiple solutions by caching the feasible solutions found during the search process or modifying hyperparameters such as budgets or seed solutions. These methods can guarantee feasibility and optimality; however, they cannot synthesize insights from data such as historical good design examples or generate unstructured objects like images that are difficult to handle explicitly in optimization problems.
Infeasibility penalization
Further work has endeavored to modify deep generative models such as generative adversarial networks (Goodfellow et al., 2014) to penalize constraint violation. General purpose methods are proposed for generating objects respecting constraint graphs (Para et al., 2021) and blackbox constraints (Di Liello et al., 2020). Specialized approaches consider graph constraints where the generated object should be a graph that meets certain constraints that are specified by a blackbox function such as in the design of photonic crystals (Christensen et al., 2020), crystal structure prediction (Kim et al., 2020), graph-constrained house generation (Tang et al., 2023), and general house plan generation (Nauata et al., 2021). In these settings, generating infeasible objects is penalized, and the generative model often has an inductive bias in the network architecture to generate feasible objects. These methods can use historical training datasets and handle unstructured objects; however, the output is often infeasible.
Optimization-based priors
In a related direction, previous work proposed conditioning variational autoencoders (VAE) (Kingma & Welling, 2013) with combinatorial programs (Misino et al., 2022). Here, the latent information is extracted by the encoder and then fed through a logic program, made differentiable via DeepProblog (Manhaeve et al., 2018), such as performing handwritten operations on digits. The result is then fed through the decoder to generate the original image based on logical relationships. At test time, the goal is to generate objects that hopefully satisfy the logical relationship because the model was conditioned to do so. Here, the logical program helps to condition the generative model to generate objects based on patterns resulting from logical relationships and penalize cases where the logic is incorrect. Here, the generated objects are not guaranteed to satisfy the combinatorial constraints, but rather, the generative model has a logical structural prior. Previous work has made various optimization problems differentiable, such as quadratic programs (Amos & Kolter, 2017), probabilistic logic programs (Manhaeve et al., 2018), linear programs (Wilder et al., 2019a; Mandi & Guns, 2020; Elmachtoub & Grigas, 2017; Liu & Grigas, 2021), Stackelberg games (Perrault et al., 2020), normal form games (Ling et al., 2018), kmeans clustering (Wilder et al., 2019b), maximum likelihood computation (Niepert et al., 2021), graph matching (Rolínek et al., 2020), knapsack (Demirovic et al., 2019b; a), maxsat (Wang et al., 2019), mixed integer linear programs (Mandi et al., 2020; Paulus et al., 2021; Ferber et al., 2020), blackbox combinatorial solvers (Pogančić et al., 2020; Mandi et al., 2022; Berthet et al., 2020), nonlinear programs (Donti et al., 2017), continuous constraint satisfaction (Donti et al., 2020), cone programs (Agrawal et al., 2019). General-purpose methods are presented for minimizing downstream regret by learning surrogate loss functions (Shah et al., 2022; 2023; Zharmagambetov et al., 2023). Additionally, previous work has employed learnable linear solvers to solve nonlinear combinatorial problems (Ferber et al., 2023). Finally, a survey investigates the intersection between machine learning and optimization (Kotary et al., 2021). These approaches focus on identifying a single solution rather than generating diverse solutions. Additionally, many of these approaches are amenable for use by GenCO if a specific optimization problem better suits the generative problem.
Enhacing combinatorial optimization with generative models
Recently, generative models have been proposed to improve combinatorial optimization. Zhang et al. (2022) use generative flow networks (gflownet) for robust scheduling problems. Sun & Yang (2023) use graph diffusion to solve combinatorial problems on graphs. Ozair et al. (2021) use vector quantized variational autoencoders to compress the latent space for solving planning problems in reinforcement learning. Zhao & You (2020) use generative adversarial networks to generate settings for sample average approximation in robust chance-constrained optimization. Additionally, in (Lopez-Piqueres et al., 2023), the authors generate continuous objects with tensor networks that satisfy linear constraints for optimization problems. These approaches target solving a fully specified optimization problem rather than generating objects with combinatorial constraints in mind.
3 GenCO: Method Description
Figure 1: An overview of GenCO. The top part of the diagram corresponds to the ordinary “GAN-style” generative model (chosen as an example). By integrating the “combinatorial loss” into both the loss function and the object generation process (bottom part), we effectively encode the combinatorial nature of the problem.
3.1 Mathematical Formulation
The key distinction in our framework centers around the incorporation of a combinatorial loss, a departure from conventional deep generative models like Generative Adversarial Networks (GANs) illustrated in Figure 1. Rather than directly generating the ultimate object of interest, be it images or any other complex entity, our generator $G{}(\epsilon{};\theta{})$ parameterized by $\theta{}$ takes random noise $\epsilon{}$ and transforms it into a problem representation $c$ that encapsulates the underlying problem’s essential features. This representation serves as a pivotal intermediary step. Subsequently, we feed $c$ to both the combinatorial loss and the generator loss (e.g. derived via a discriminator), ensuring that the solutions produced by the generator exhibit the desired characteristics: diversity, realism, and adherence to combinatorial properties.
To train such a generator, we define the following bi-objective function, effectively balancing the generative loss $\mathcal{L}_{\text{gen}}{}$ with the combinatorial loss $\mathcal{C}{}$:
$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(c)+\gamma\mathcal{C}{}(h(c))]$ (1)
s.t. $c=G(\epsilon;\theta{}),$
where $\mathcal{L}_{\text{gen}}{}$ encapsulates the generative loss, responsible for promoting diversity and realism. Its form can vary substantially depending on the specific generative model employed. For example, a Wasserstein GAN might utilize the adversarial discrepancy between outputs, while a Variational Autoencoder (VAE) would make use of the Evidence Lower Bound (ELBO).
Conversely, the combinatorial loss $\mathcal{C}{}(\cdot)$, penalized by the weighting factor $\gamma$, is coupled with the downstream optimization solver $h(\cdot)$. This solver plays a crucial role in enforcing or encouraging combinatorial properties. For instance, one might seek to generate objects that must adhere to physical or logical constraints, all while minimizing a specified (linear) cost. In such a scenario, it becomes imperative to formulate an instance of a Mixed-Integer Linear Program (MILP) or another type of optimization problem with a well-defined feasible region. This problem instance then integrates seamlessly into the training pipeline, serving as the backbone of $\mathcal{C}{}$.
Below, we consider two potential and practical settings for $\mathcal{C}{}$, each tailored to address specific real-world applications.
3.2 Constrained Generator
In this setting, we consider a problem domain where the goal is to train a generator that outputs feasible solutions, which natively handles diversity. To generate objects that are guaranteed to be feasible, we set $\gamma\to+\infty$ in equation 1 for any $c$ that does not satisfy constraints (defined by feasible region $\Omega$), 0 otherwise. To handle this directly with such a high value of $\gamma$ is computationally inefficient. Instead, we employ a ”projection layer” denoted as $h(\cdot)$ that maps $c$ to its nearest feasible point in the feasible set $\Omega$. Specifically, we define $h(c)=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$  [^1] Although a dot product is a natural choice of proximity for cosine distance, one could employ other metrics, e.g. $|c-x|$, where $c$ is an output of the generator $G(\epsilon;\theta)$. In this context, a generated problem description $c$ (as in Fig.1) can be viewed as an unconstrained generated object that is subsequently mapped via $h(\cdot)$ to $x$ to adhere to the defined constraints. With this approach, we can reformulate equation 1 into the following unconstrained optimization problem:
$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(h(G(\epsilon;\theta{})))].$ (2)
Note that the feasibility within region $\Omega$ is guaranteed to be satisfied due to the form of $h(\cdot)$. Directly minimizing this objective is possible thanks to the recent advancements in differentiable solvers literature (Agrawal et al., 2019; Pogančić et al., 2020; Sahoo et al., 2022).
Algorithm 1 provides a high-level description of our method for this setting. It starts by initializing the generator parameters. In each iteration, a problem description is randomly selected using the generator. An optimization problem is then solved to find the solution that maximizes the similarity between $c$ within a defined feasible set ($x$ is the corrected/feasible version of $c$). Finally, the generator parameters are updated through backpropagation. We also provide an application of this generic algorithm to the concrete generative model examples of GANs and VQVAEs in Appendices B-C.
  Output: Trained generator
  Initialize generator parameters $\theta$;
  while not converged do
     Sample a noise $\epsilon$;
     Sample a problem description $c\sim G(\epsilon;\theta)$;
     Call a solver $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x=h(c)$;
     Compute $\text{loss}=\mathcal{L}_{\text{gen}}{}(x^{*})$;
     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$;
  end while
Algorithm 1 Generator Training in Constrained Setting
3.3 Penalized Generator
The formulation from the previous section 3.2 naturally extends to penalty form, i.e., we encourage generated problem representation $c$ to respect certain combinatorial properties. In this scenario, we can directly leverage equation 1 as there is no need in projection layer:
$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(G(\epsilon;\theta{}))+\gamma\mathcal{C}{}(h(G(\epsilon;\theta{})))].$ (3)
Here, we seamlessly integrate the generator into the objective function. While it is true that we do not impose any explicit constraints on the output of the generator, and one could use the output of $G$ directly, we contend that introducing a combinatorial penalty or loss as a penalty remains a crucial framework with significant practical utility. This approach serves to further guide the generation process and injects domain-specific knowledge that is easy to interpret, ensuring that the generated solutions exhibit desired combinatorial properties, a facet that holds immense practical relevance.
For instance, we may aim to generate images tailored for path planning in strategy games. These maps feature various types of terrains, each incurring distinct costs when placed on the map. For efficient path planning, our goal is to find a “low cost” shortest path between two endpoints. Therefore, in this scenario, $\mathcal{C}$ incorporates a domain specific (fixed) mapping (e.g. a neural net) that transforms an image into a graph with corresponding weights assigned to each edge, followed by solving the shortest path problem (e.g. using Dijkstra’s algorithm).
However, in this particular case, there is no need to introduce a “projection” operator as in the previous case since there are no explicit constraints that must be strictly satisfied. We can directly employ gradient-based optimization, allowing for differentiation through the combinatorial problem. This is achieved approximately by employing methods outlined in (Pogančić et al., 2020; Sahoo et al., 2022). Pseudocode 2 describes the application of GenCO for penalty setting. Furthermore, Appendix A.1 provides an application of this generic algorithm to a concrete example of GANs.
  Output: Trained generator
  Initialize generator parameters $\theta$;
  (optional) Define a fixed mapping $g(\cdot)$ for $c$
  while not converged do
     Sample a noise $\epsilon$;
     Sample a problem description $c\sim G(\epsilon;\theta)$;
     (optional) apply a mapping $g(\cdot)$ to transform $c$;
     Call a solver $x^{*}=h(g(c))$ ;
     Compute $\text{loss}=\mathcal{L}_{\text{gen}}{}(c)+\gamma\mathcal{C}{}(x^{*})$;
     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$;
  end while
Algorithm 2 Generator Training in Penalty Setting
It is crucial to emphasize that the given penalized setting deviates from the conventional decision-focused learning (DFL) or smart predict-then-optimize paradigms (Elmachtoub & Grigas, 2017; Donti et al., 2017). In our approach, we do not possess access to the ground truth costs $c$, and the ultimate loss function encompasses both the generator’s loss and the output of a solver.
4 Experiments
We tested our approach on two applications: level generation for the Zelda game and map generation for path planning in Warcraft. Although both applications are in the game domain, they have quite different use cases. Nevertheless, both settings involve combinatorial optimization in the pipeline, which makes the typical use of deep generative models nontrivial. In these experiments, we demonstrated that GenCO significantly outperforms the baselines, efficiently finding diverse, realistic solutions that obey combinatorial properties. This success paves the way for combining combinatorial optimization with deep generative models.
4.1 Generating game levels using GANs
We evaluate our model on the task of generating diverse Zelda game levels. In this setting, we use GenCO to train using a GAN-like constrained formulation (refer to section 3.2 and Appendix B).
4.1.1 Settings
Explicitly Constrained GAN: Game Level Design
We train GenCO on the task of generating Zelda game levels. Here, we are given examples of human-crafted game levels and are asked to generate fun new levels. The generated levels must be playable in that the player must be able to solve them by moving the character through a route that reaches the destination. Additionally, the levels should be realistic in that they should be somewhat similar to the real game levels as measured by an adversary that is trained to distinguish between real and fake images ($\mathcal{L}_{\text{gen}}{}$). We use the same dataset as (Zhang et al., 2020), consisting of 50 Zelda game levels, as well as the same network architectures, as we don’t tune the hyperparameters for our model in particular but rather compare the different approaches, all else equal.
Here, we evaluate several approaches, including the previous work (Zhang et al., 2020), which we call GAN + MILP fix. This approach first trains a standard Wasserstein GAN architecture to generate the game levels. Specifically, they train the WGAN by alternatively training two components: a generator and a discriminator. The generator is updated to “fool” the discriminator as much as possible in that it tries to maximize the loss of the discriminator. The discriminator tries to correctly separate the generated and real game levels into their respective classes. When the practitioner then wants to generate a valid level, this approach generates a level with the generator and then fixes it using a MILP formulation that finds the nearest feasible game level where proximity is determined by cosine distance, which is equivalent to minimizing a dot product distance.
We evaluate two variants of GenCO, GenCO - Fixed Adversary, which iteratively updates the generative model to fool a fixed pretrained adversary, and GenCO - Updated Adversary, which updates both the generator and adversary during training. The fixed adversary approach trains to fool a pretrained adversary that is obtained from the fully trained GAN from previous work (Zhang et al., 2020). Both GenCO approaches are initialized with the fully trained GAN from previous work (Zhang et al., 2020).
4.1.2 Results
Approach % unique
GAN loss ($\mathcal{L}_{\text{gen}}{}$)
(lower better)
GenCO adversary
(lower better)
GenCO - Updated Adversary 0.995 -10.10 -4.49
GenCO - Fixed Adversary 0.22 -1.45 -0.85
GAN + MILP fix (Previous Work) 0.52 0.22 0.24
Table 1: Game level design comparison. We compare GenCO with an updating adversary against GenCO with a fixed adversary, and lastly, previous work that postprocesses solutions to be valid game levels. Here, all levels are valid, but GenCO exhibits more diversity and realism.
GAN + MILP Fix Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO Fixed Adv Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO Updated Adv Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Generated zelda game examples. The GenCO Updated Adversary obtains solutions that seem more realistic than mainly empty fixed GAN instances and fixed adversary instances.
We present a table of results in Table 1 and examples in Figure 2, which includes the performance of the previous approach as well as two variants of our proposed approach. In these settings, we estimate performance based on sampling 1000 levels. Each level is made out of a grid, with each grid cell having one of 8 components: wall, empty, key, exit, 3 enemy types, and the player. A valid level is one that can be solved by the player in that there is a valid route starting at the player’s location, collecting the key, and then reaching the exit. We evaluate the performance of the models using two types of metrics: diversity, as measured by the percentage of unique levels generated, and fidelity, as measured by the average objective quality of a fixed GAN adversary. We evaluate using adversaries from both the previous work and GenCO. As shown in Table 1, we find that GenCO with an updating adversary generates unique solutions at a much higher rate than previous approaches and also generates solutions that are of higher quality as measured by both the GAN adversary and its own adversary. The adversary quality demonstrates that the solutions are realistic in that neither the adversary from the previous work nor from GenCO is able to distinguish the generated examples from the real examples. This is further demonstrated in Figure 2 with the updated adversary generating realistic and nontrivial game levels. Furthermore, given that the levels are trained on only 50 examples, we can obtain many more game levels. Note here that both approaches are guaranteed to give playable levels as they are postprocessed to be valid. However, GenCO is able to generate more diverse solutions that are also of higher quality.
Uniqueness
As shown in Table 1, GenCO, with an updated adversary, obtains the highest percentage of unique solutions, generating 995 unique solutions out of 1000. This is significantly higher than the previous work, which only generated 520 unique solutions. This is likely due to the fact that the generator is trained with the downstream fixing explicitly in the loop. This means that while the previous work may have been able to “hide” from the adversary by generating slightly different continuous solutions, these continuous solutions may project to the same discrete solution. On the other hand, by integrating the fixing into the training loop, GenCO’s generator is unable to hide in the continuous solution space and thus is heavily penalized by generating the same solution as the adversary will easily detect those to be originating from the generator. In essence, this makes the adversary’s task easier as it only needs to consider distinguishing between valid discrete levels rather than continuous and unconstrained levels. This is also reflected in the adversary quality, where GenCO’s adversary is able to distinguish between levels coming from the previous work’s generator and the real levels with a much better loss.
4.2 Map generation for path planning
In this experiment, we consider somewhat different and more challenging for GenCO setting of generation of image maps for strategy games like Warcraft. This poses a greater challenge because the generator $G$ described in equation 1 is not restricted and can produce any image of a map (setting descrbied in section 3.3). However, we are required to encourage $G$ to generate a map with two main criteria: 1) it should resemble real and diverse game maps, as determined by an adversary trained to differentiate between real and fake images; 2) the cost of the shortest path from the top-left to the bottom-right corners (source and destination) should be minimized. Intuitively, it corresponds to (mostly) a diagonal part of the image, and a map can include various elements (terrains) like mountains, lakes, forests, and land, each with a specific cost (for example, mountains may have a cost of 3, while land has a cost of 0). To calculate this, we pass the generated image to the fixed ResNet (called “cost NN”) to get the graph representation together with edge costs. The cost NN corresponds to the mapping $g(\cdot)$ in Algorithm 2. This is then followed by the shortest path solver ($h(\cdot)$ in Algorithm 2). The objective is to populate the map with a diverse range of objects while ensuring the shortest path remains low-cost. We use the same dataset as in (Pogančić et al., 2020) with DCGAN architecture adapted from (Zhang et al., 2020). More implementation details and experimental settings can be found in Appendix A.
4.2.1 Baselines
We examine various baseline models, including an “Ordinary GAN” that does not incorporate the Shortest Path objective (see the top part of Fig. 1). Like the Zelda experiment, this approach employs a standard Wasserstein GAN architecture closely following (Zhang et al., 2020). However, in this case, we generate images directly instead of encoding game levels. More precisely, we train the WGAN by iteratively training two components: a generator and a discriminator. The generator is fine-tuned to deceive the discriminator to the greatest extent possible, aiming to maximize the discriminator’s loss. The discriminator, on the other hand, endeavors to accurately distinguish between generated and authentic game levels and categorize them accordingly.
The next baseline is “GAN + cost NN”, which simultaneously feeds the generator’s output into both the discriminator and a ResNet that calculates the costs for each edge in the grid. We then average the output to obtain the final loss for generator. We then average these outputs to derive the ultimate loss for the generator. While this approach considers costs associated with objects, it does not incorporate information about the Shortest Path.
As for the GenCO, it generalizes both of these approaches incorporating combinatorial solver into the pipeline. The detailed algorithm is given in Appendix A. To backpropagate through the solver, we employ the “identity with projection” method from (Sahoo et al., 2022). The remaining settings are similar to “Ordinary GAN”.
4.2.2 Results
Quantitative results are showcased in Table 2. The “Ordinary GAN” focuses exclusively on the GAN’s objective, without considering the Shortest Path’s objective $f$. In contrast, GenCO achieves higher performance with regards to the SP’s objective, albeit with a slight reduction in GAN’s loss. On the other hand, the “GAN + cost NN” approach optimizes using the cost vector of the entire map (generated by ResNet). In other words, it tries to uniformly avoid placing costly objects in any part of the image, whereas we are interested only in the shortest path. While it demonstrates a modest enhancement in the objective $f$, it experiences a notable decline in terms of generator loss ($\mathcal{L}_{\text{gen}}$). This indicates a trade-off between optimizing for the Shortest Path and the GAN’s objective.
Approach
GAN loss ($\mathcal{L}_{\text{gen}}$)
(lower better)
SP loss ($\mathcal{C}$)
(lower better)
Ordinary GAN 0.6147 36.45
GAN + cost NN 0.8994 35.61
GenCO (ours) 0.6360 23.99
Table 2: Performance comparison on Warcraft map generation. Our here goal is to create a map that is both realistic (“fools” discriminator, i.e. lower GAN loss) and has small objective for the shortest path (SP). Results are averaged across 100 instances.
Such quantitative results directly translate into image qualities. In fig. 3, a subset of generated Warcraft map images using various methods is displayed. The “Ordinary GAN” tends to generate maps with elements such as mountains and lakes, which are considered “very costly”, particularly along the Shortest Path from the top-left to the bottom-right (which mostly goes through the diagonal). This makes sense since WGAN is trained with no information about grid costs. In contrast, the “GAN + cost NN” approach produces less costly maps, albeit with reduced diversity. For instance, most part of the image is populated with the same object. On the other hand, GenCO strikes a balance, achieving a cheap Shortest Path while maintaining a diverse range of elements on the map. The Shortest Path is efficient, and the map exhibits a rich variety of features simultaneously.
Ordinary GAN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GAN + cost NN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO (ours) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: A subset of generated Warcraft map images using various methods. “Ordinary GAN” generates “very costly” maps (e.g. mountains, lakes) along the Shortest Path from top-left to bottom-right; “GAN + cost NN” generates less costly maps but they are less diverse; GenCO: SP path is cheap and the map is diverse at the same time.
5 Conclusion
In this paper, we introduce GenCO, a framework for integrating combinatorial constraints in a variety of generative models, and show how it can be used to generate diverse combinatorial solutions to nonlinear optimization problems. Unlike existing generative models and optimization solvers, GenCO guarantees that the generated diverse solutions satisfy combinatorial constraints, and we show empirically that it can optimize nonlinear objectives.
The underlying idea of our framework is to combine the flexibility of deep generative models with the guarantees of optimization solvers. By training the generator end-to-end with a surrogate linear combinatorial solver, GenCO generates diverse and combinatorially feasible solutions, with the generative loss being computed only on feasible solutions.
We have tested GenCO on various combinatorial optimization problems and generative settings, including GAN in Zelda game level generation and Warcraft map generation for path planning, demonstrating the flexibility of our framework for integrating into different generative paradigms. Our framework consistently produced diverse and high-quality solutions that satisfy the combinatorial constraints, which can be flexibly encoded using general-purpose combinatorial solvers.
Appendix A Detailed experimental setup for map generation
A.1 Pseudocode
1:  Initialize generator parameters $\theta_{gen}$;
2:  Initialize adversary parameters $\theta_{adv}$;
3:  Input: distribution of the true dataset ($c_{true}\sim p_{\text{data}(c)}$), GAN’s objective $\mathcal{L}(\theta_{gen},\theta_{adv})$;
4:  for epoch $e$ do
5:     for batch $b$ do
6:        Sample a noise $\epsilon$;
7:        Sample true examples from dataset $c_{true}\sim p_{\text{data}}(c)$;
8:        Sample fake examples using $c_{fake}\sim G(\epsilon;\theta_{gen})$;
9:        Transform $c_{fake}$ into the coefficients of the optimization problem: $c=g(c_{fake})$;
10:        Solve: $x^{*}=\operatorname*{arg\,min}_{x\in\Omega}c^{T}z$;
11:        Backpropagate $\nabla_{\theta_{gen}}\left[\mathcal{L}(c_{fake};\theta_{gen},\theta_{adv})+\beta c^{T}x^{*}\right]$ to update $\theta_{gen}$;
12:        Backpropagate $\nabla_{\theta_{adv}}\left[-\mathcal{L}(c_{fake};\theta_{gen},\theta_{adv})+\mathcal{L}(c_{real};\theta_{gen},\theta_{adv})\right]$ to update $\theta_{adv}$;
13:     end for
14:  end for
Algorithm 3 GenCO Penalized generator training for GANs.
Algorithm 3 provides a detailed description of the GenCO framework in our penalty formulation and utilized in section 4.2. In this process, we sample both real and synthetic data, drawing from the true data distribution and the generator $G$ respectively (lines 7–8). Subsequently, the synthetic data undergoes a fixed mapping (e.g. ResNet in our experiments), called cost neural net (or cost NN), to obtain coefficients for the optimization problem, specifically the edge weights for the Shortest Path. Following this, we invoke a solver that provides us with a solution and its associated objective (lines 10–11). We then proceed to update the parameters of the generator $G$ using both the GAN’s objective and the solver’s objective. Finally, we refine the parameters of the adversary (discriminator) in accordance with the standard GAN’s objective.
A.2 Settings
We employ ResNet as the mapping $g(\cdot)$ from Algorithm 3, which transforms an image of a map into a $12\times 12$ grid representation of a weighted directed graph: $g:\Re^{96\times 96\times 3}\rightarrow\Re^{12\times 12}$. The first five layers of ResNet18 are pre-trained (75 epochs, Adam optimizer with lr=$5e-4$) using the dataset from Pogančić et al. (2020), comprising 10,000 labeled pairs of image–grid (refer to the dataset description below). Following pretraining, we feed the output into the Shortest Path solver, using the top-left point as the source and the bottom-right point as the destination. The resulting objective value from the Shortest Path corresponds to $f$.
Dataset
The dataset used for training in the Shortest Path problem with $k=12$ comprises 10,000 randomly generated terrain maps from the Warcraft II tileset Pogančić et al. (2020) (adapted from Guyomarch (2017)). These maps are represented on a $12\times 12$ grid, with each vertex denoting a terrain type and its associated fixed cost. For example, a mountain terrain may have a cost of 9, while a forest is assigned a cost of 1. It’s important to note that in the execution of Algorithm 3, we don’t directly utilize the actual (ground truth) costs, but rather rely on ResNet to generate them.
A.3 Architecture
We employ similar DCGAN architecture taken from Zhang et al. (2020) (see fig. 3 therein). Input to generator is 128 dimensional vector sampled from Gaussian noise centered around 0 and with a std of 1. Generator consists of five (256–128–64–32–16) blocks of transposed convolutional layers, each with $3\times 3$ kernel sizes and batch normalization layers in between. Discriminator follows by mirroring the same architecture in reverse fashion. The discriminator mirrors this architecture in reverse order. The entire structure is trained using the WGAN algorithm, as described inZhang et al. (2020).
Appendix B GANs with combinatorial constraints
In the generative adversarial networks (GAN) setting, the generative objectives are measured by the quality of a worst-case adversary, which is trained to distinguish between the generator’s output and the true data distribution. Here, we use the combinatorial solver to ensure that the generator’s output is always feasible and that the adversary’s loss is evaluated using only feasible solutions. This not only ensures that the pipeline is more aligned with the real-world deployment but also that the discriminator doesn’t have to dedicate model capacity to detecting infeasibility as indicating a solution is fake and instead dedicate model capacity to distinguishing between real and fake inputs, assuming they are all valid. Furthermore, we can ensure that the objective function is optimized by penalizing the generator based on the generated solutions’ objective values:
$\mathcal{L}_{\text{gen}}{}(G(\epsilon;\theta_{gen}))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(G(\epsilon;\theta_{gen})))\right]$ (4)
where $f_{\theta_{adv}}$ is an adversary (a.k.a. discriminator) and putting this in the context of equation 2 leads to:
$\min_{\theta_{gen}}\quad$ $\mathcal{L}_{\text{gen}}{}(S(\tilde{G}(\epsilon;\theta_{gen})))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(S(\tilde{G}{}(\epsilon;\theta_{gen}))))\right]$ (5)
where $\tilde{G}{}$ is unconstrained generator and $S$ is a surrogate combinatorial solver as described above. Here, we also have adversary’s learnable parameters $\theta_{adv}$. However, that part does not depend on combinatorial solver and can be trained as in usual GAN’s. The algorithm is presented in pseudocode 4.
  Initialize generator parameters $\theta_{gen}$;
  Initialize adversary parameters $\theta_{adv}$;
  for epoch $e$ do
     for batch $b$ do
        Sample problem $\epsilon{}$;
        Sample true examples from dataset $x_{true}\sim p_{\text{data}}(x)$;
        Sample linear coefficients $c\sim G(\epsilon{};\theta_{gen})$;
        Solve $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$;
        Backpropagate $\nabla_{\theta_{gen}}\mathcal{L}_{\text{gen}}{}(x^{*};\theta_{gen})$ to update generator (equation 5);
        Backpropagate $\nabla_{\theta_{adv}}\left[\log(f_{\theta_{adv}}(x_{true}))-\mathcal{L}_{\text{gen}}{}(x^{*};\theta_{gen})\right]$ to update adversary;
     end for
  end for
Algorithm 4 GenCO GenCO in the constrained generator setting
Appendix C GenCO – VQVAE
The formulation below spells out the VQVAE training procedure. Here, we simply train VQVAE on a dataset of known objective coefficients, which solves the problem at hand. A variant of this also puts the decision-focused loss on the generated objective coefficients, running optimizer $g_{\text{solver}}$ on the objective coefficients to get a solution and then computing the objective value of the solution.
$L_{\text{ELBO}}(c,\theta,E)=\mathbb{E}_{q_{\theta}(z|c)}[\log p_{\theta}(c|z)]-\beta\cdot D_{\text{KL}}(q_{\theta}(z|c)||p(z))+\gamma\cdot\|sg(\mathbf{e}_{k})-\mathbf{z}_{e,\theta}\|_{2}^{2}$ (6)
Here $z$ is an embedding vector, $c$ is the objective coefficients, $\log p_{\theta}(c|z)$ is a loss calculated via the mean squared error between the decoder output and the original input objective coefficients, $q_{\theta}(z|c)$ is the encoder, $p(z)$ is the prior, and $sg(\cdot)$ is the stop gradient operator, $E$ is a discrete codebook that is used to quantize the embedding.
$\mathcal{L}_{\text{optimization}}=\mathbb{E}_{c\sim p_{\theta}(c\mid z)}\left[f_{\text{obj}}(g_{\text{solver}}(c;y))\right]$ (7)
The algorithm below maximizes a combination of the losses in Equation equation 6 and Equation equation 7.
Input:Training data distribution $D$ over problem info $y$ and known high-quality objective coefficients $c$, regularization weight $\beta$, linear surrogate solver $g_{\text{solver}}$, nonlinear objective $f_{\text{objective}}$. Output:Trained encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and codebook $E$ Initialize the parameters of the encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and the codebook $E=\{e_{1},e_{2},\ldots,e_{K}\}$ with $K$ embedding vectors;
for $t=1$ to $T$ do
       Sample $y,c$ from the distribution $D$;
       Compute the encoder output $z_{e}=f_{\text{enc}}(y,c)$;
       Find the nearest embedding vector $z_{q}=\arg\min_{e\in E}\|z_{e}-e\|_{2}^{2}$;
       Compute the quantization loss $L_{\text{quant}}=\|z_{e}-z_{q}\|_{2}^{2}$;
       Compute the reconstruction $\tilde{c}=f_{\text{dec}}(y,z_{q})$;
       Compute the reconstruction loss $L_{\text{recon}}=\|c-\tilde{c}\|_{2}^{2}$;
       GenCo Variant: Compute the optimization loss $L_{\text{opt}}=f_{\text{objective}}(g_{\text{solver}}(\tilde{c},y))$;
       Compute the total loss: $L_{\text{total}}=L_{\text{recon}}+\beta_{1}L_{\text{quant}}$ + $\beta_{2}L_{\text{opt}}$;
       Update the parameters of the encoder, decoder, and codebook to minimize $L_{\text{total}}$;
end for
Algorithm 5 Constrained generator Training for VQVAE

### v2

@@
-GenCO: Generating Diverse Solutions to Design Problems with Combinatorial Nature
+GenCO: Generating Diverse Designs with Combinatorial Constraints
 Abstract
-Generating diverse objects (e.g., images) using generative models (such as GAN or VAE) has achieved impressive results in the recent years, to help solve many design problems that are traditionally done by humans.
-Going beyond image generation, we aim to find solutions to more general design problems, in which both the diversity of the design and conformity of constraints are important.
-Such a setting has applications in computer graphics, animation, industrial design, material science, etc, in which we may want the output of the generator to follow discrete/combinatorial constraints and penalize any deviation, which is non-trivial with existing generative models and optimization solvers.
-To address this, we propose GenCO, a novel framework that conducts end-to-end training of deep generative models integrated with embedded combinatorial solvers, aiming to uncover high-quality solutions aligned with nonlinear objectives.
-While structurally akin to conventional generative models, GenCO diverges in its role - it focuses on generating instances of combinatorial optimization problems rather than final objects (e.g., images).
-This shift allows finer control over the generated outputs, enabling assessments of their feasibility and introducing an additional combinatorial loss component.
-We demonstrate the effectiveness of our approach on a variety of generative tasks characterized by combinatorial intricacies, including game level generation and map creation for path planning, consistently demonstrating its capability to yield diverse, high-quality solutions that reliably adhere to user-specified combinatorial properties.
+Deep generative models like GAN and VAE have shown impressive results in generating unconstrained objects like images.
+However, many design settings arising in industrial design, material science, computer graphics and more require that the generated objects satisfy hard combinatorial constraints or meet objectives in addition to modeling a data distribution.
+To address this, we propose GenCO, a generative framework that guarantees constraint satisfaction throughout training by leveraging differentiable combinatorial solvers to enforce feasibility.
+GenCO imposes the generative loss on provably feasible solutions rather than intermediate soft solutions, meaning that the deep generative network can focus on ensuring the generated objects match the data distribution without having to also capture feasibility.
+This shift enables practitioners to enforce hard constraints on the generated outputs during end-to-end training, enabling assessments of their feasibility and introducing additional combinatorial loss components to deep generative training.
+We demonstrate the effectiveness of our approach on a variety of generative combinatorial tasks, including game level generation, map creation for path planning, and photonic device design, consistently demonstrating its capability to yield diverse, high-quality solutions that verifiably adhere to user-specified combinatorial properties.
+Machine Learning, ICML
+[…]
+usc,cornell,FAIR
 1 Introduction
 Generating diverse and realistic objects with combinatorial properties is an important task with many applications.
-For example, in video game level design, we may want to generate a variety of levels that are both realistic and valid/playable (Zhang et al., 2020).
-Here, “valid” may refer to certain discrete characteristics of the level that must be satisfied (e.g., a minimum number of enemies, a path between the level entrance and exit, etc.).
-In automatic device design, we want to generate a variety of devices that meet foundry manufacturing constraints and optimize physics-related objectives (Schubert et al., 2022).
+For AI-guided design in particular, the goal is to generate a diverse set of combinatorial solutions that are both feasible and high quality with respect to a given nonlinear objective.
+Diversity is often useful when dealing with uncertain metrics, diverse preferences, or for the sake of further manual scrutiny.
+For example, in photonic device design, we want to generate a variety of devices that meet foundry manufacturing constraints and optimize physics-related objectives (Schubert et al., 2022).
 In the design of new molecules, we want to generate a variety of chemically valid molecules with specific properties (Pereira et al., 2021).
-In all of these examples, the goal is to generate a variety of combinatorial solutions that are both feasible and high quality with respect to a given nonlinear objective.
-Another situation to consider is when the generated solutions do not have to strictly follow combinatorial constraints, but we want to discourage certain things as a penalty.
-For example, if we are creating various images of a map, we might want to ensure that the path planning in those maps is efficient (see section 4.2).
-The difficulty in these settings arises from the fact that the generated objects need to satisfy specific discrete properties incorporated either as combinatorial constraints or penalties and be of high quality as evaluated by a nonlinear objective.
-The combinatorial constraints make it challenging to apply gradient-based methods to train standard generative models, such as GAN (Goodfellow et al., 2014), or VAE (Kingma & Welling, 2013), which may generate examples that violate the user-specified constraints.
-Additionally, the nonlinear objective makes it difficult to use traditional general-purpose combinatorial optimization tools since mixed-integer nonlinear programming (MINLP) is often slow in practice compared to the more heavily researched mixed integer linear programming (MILP), often requiring problem-specific solvers or leveraging specialized structure.
-Furthermore, traditional optimization solvers are not geared towards generating diverse solutions and do not have the flexibility of deep learning models to quickly adapt to slightly modified settings.
-Recent work has attempted to address these issues by first training a generative model and then fixing/postprocessing the output after the fact (Zhang et al., 2020).
-However, these approaches are limited in that they only use the combinatorial solver to fix the output of the generative model and do not use the combinatorial solver in training.
-As a result, the generative model is not trained with the combinatorial solver in mind (i.e., end-to-end).
-Thus, the generator may fail to capture the distribution of high-quality feasible solutions as many of the generated examples may be “fixed” to a single solution.
-Additionally, some work constrains generative model outputs by penalizing the generator based on constraint violation (Chao et al., 2021).
-While this approach may work in some settings where the constraints must be satisfied, it is unclear how to extend this approach to more complex constraints, such as logical or general combinatorial constraints that can be readily expressed in MILP.
-As a result, we propose GenCO that generates diverse solutions to design problems with combinatorial nature.
-GenCO combines the flexibility of deep generative models with the combinatorial efficiency of optimization solvers.
-Our approach mainly deals with two cases: (1) when the resulting solution needs to satisfy certain hard combinatorial constraints (hard constraints), and (2) the imposition of a desired penalty if the solutions violate the constraints (soft constraints).
-To achieve these two goals, we introduce a general concept known as the “combinatorial loss”, which serves as an additional “restriction” on the generated solutions.
-In the context of combinatorial constraints, the combinatorial loss takes the form of a projection expressed as a mixed-integer linear program (MILP).
-In the second scenario, this general concept manifests as an extra penalty term within the objective function.
-This, coupled with the diversity mechanisms inherent in deep generative models, fine-tunes the generated solutions.
-This process both encourages diversity and ensures adherence to the specified combinatorial properties.
-Our main contributions involve introducing a framework, GenCO, that serves two key purposes: 1) generating a range of solutions to nonlinear problems, ensuring they adhere to combinatorial properties; 2) integrating a combinatorial solver into the deep generative learning process.
-We showcase the effectiveness of our approach through a series of experiments on various generative combinatorial optimization problems.
-This includes tasks like game level design (section 4.1) and generating maps for path planning (section 4.2).
+In other scenarios, we would like to create a diverse set of designs so that their solutions (not necessarily diverse given the design) meet certain objectives.
+For example, in video game level design, we may want to generate levels that are both realistic and valid/playable (Zhang et al., 2020).
+Here, “valid” may refer to discrete characteristics of the level that must be satisfied (e.g., a minimum number of enemies, a guaranteed path between the level entrance and exit, etc.).
+Approaches Solution Feasibility Data Nonlinear
+generation guarantees distribution objective
+GAN/VAE (Goodfellow et al., 2014; Kingma & Welling, 2013) Fast
+Integer Programming (Deaton & Grandhi, 2014) Slow
+Lagrangian Multiplier (Di Liello et al., 2020) Slow
+SurCo (Ferber et al., 2023) Slow
+Fixing/Rejection (Zhang et al., 2020; Gomes et al., 2006) Slow
+GenCO (proposed) Fast
+Table 1: A comparison of relevant directions tackling generative modeling or combinatorial optimization for design problems.
+Mathematically, we seek several solutions $\mathcal{X}:=\{{\bm{x}}_{j}\}$ to the following optimization problem:
+$\min_{\mathcal{X}}\mathcal{L}(\mathcal{X})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})\quad\quad\mathrm{s.t.}\ {\bm{x}}_{j}\in\Omega$ (1)
+where $\mathcal{L}$ is a group loss defined on a population of solutions $\mathcal{X}$ (e.g., minimizing the Wasserstein distance to a dataset), $\mathcal{D}$ is an individual loss measuring the quality of each solution ${\bm{x}}_{j}$, $\gamma$ is the weighing coefficient and $\Omega$ is the feasible set.
+Eqn.
+1 can be solved by direct optimization or learning approaches that leverage solution datasets.
+Several approaches can attempt to solve this problem.
+The Lagrangian (Frangioni, 2005) can reformulate the problem as $\min_{\mathcal{X}}\mathcal{L}(\mathcal{X})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})+\lambda\sum_{j}\mathrm{dist}_{\Omega}({\bm{x}}_{j})$, where $\mathrm{dist}_{\Omega}({\bm{x}}_{j})$ measures the distance between ${\bm{x}}_{j}$ and the feasible set $\Omega$ ($\mathrm{dist}_{\Omega}({\bm{x}}_{j})=0$ if ${\bm{x}}_{j}\in\Omega$).
+However the final solutions ${\bm{x}}_{j}$ may not be feasible unless we use $\lambda\rightarrow+\infty$ which may lead to numerical instability.
+To reduce the optimization cost, we want to learn a generative model to directly generate ${\bm{x}}_{j}$, and avoid on-the-fly optimization.
+However, standard methods like GAN (Goodfellow et al., 2014) or VAE (Kingma & Welling, 2013) fail to generate feasible and high quality solutions.
+One approach (Chao et al., 2021) penalizes the generator based on constraint violation but is nontrivial to extend to more complex constraints, such as logical or general combinatorial constraints that can be readily expressed in MILP.
+Alternative approaches include postprocessing solutions (Zhang et al., 2020), which may lead to repetitive solutions since the generative network is not trained with the postprocessing, or rejection sampling (Gomes et al., 2006) that can be computationally inefficient due to high rejection rates.
+Fundamentally, the training procedure does not involve verifiably feasible solutions.
+Another approach is to leverage integer linear programming (ILP) to find feasible solutions to problems with linear constraints on integer variables.
+However, such solvers are not geared towards generating diverse solutions and cannot leverage deep learning models to capture the data distribution or quickly adapt to slightly modified settings.
+In this work, we propose GenCO.
+Instead of directly optimizing ${\bm{x}}$, we learn a latent design space $\mathcal{C}$ so that (1) each element ${\bm{c}}\in\mathcal{C}$ yields a guaranteed feasible solution ${\bm{x}}\in\Omega$, often through a combinatorial solver ${\bm{g}}$: ${\bm{x}}={\bm{g}}({\bm{c}})$, and (2) a generative model $G_{\theta}$ can map noise to the latent space $\mathcal{C}$.
+The entire pipeline is trained in an end-to-end manner via backpropagation through the solver ${\bm{g}}$.
+Intuitively, this latent space $\mathcal{C}$ represents the hidden parameters that lead to the solution.
+For example, ${\bm{c}}$ represents the layout of a 10x10 maze, while ${\bm{x}}$ represents the optimal path of the maze.
+GenCO addresses many challenges in previous work.
+GenCO generates solutions efficiently due to its learned generative model $G_{\theta}$.
+Additionally, GenCO guarantees feasibility via re-parameterization in the latent design space, which is different from conventional deep generative models like Generative Adversarial Networks (GANs) that directly generate the ultimate object of interest ${\bm{x}}$, be it images or any other complex entity.
+GenCO encourages diverse solutions, thanks to the group loss term $\mathcal{L}$ imposed on the solution population and end-to-end training of the entire pipeline.
+Finally, GenCO does not require linear objective functions and can handle nonlinear objectives.
+A comparison is summarized in  Table 1.
+Our main contributions involve introducing a framework, GenCO, that can: 1) train generative models to generate solutions that are guaranteed to satisfy combinatorial properties throughout training; 2) optimize nonlinear objectives; 3) flexibly handle a variety of generative models and optimizers.
+We showcase the effectiveness of our approach empirically on various generative combinatorial optimization problems: creating game level ( subsection 4.2), generating maps for path planning ( subsection 4.3), and designing inverse photonic devices ( subsection 4.4).
 2 Related Work
-The interaction between generative models and combinatorial optimization has seen increased research interest as practitioners seek to integrate or draw inspiration from both paradigms.
-A prominent line of research has approached the problem of generating objects with combinatorial optimization in mind.
+The interaction between generative models and combinatorial optimization has drawn attention as practitioners seek to generate objects with combinatorial optimization in mind.
 Traditional constrained object generation
@@
 Infeasibility penalization
-Further work has endeavored to modify deep generative models such as generative adversarial networks (Goodfellow et al., 2014) to penalize constraint violation.
-General purpose methods are proposed for generating objects respecting constraint graphs (Para et al., 2021) and blackbox constraints (Di Liello et al., 2020).
-Specialized approaches consider graph constraints where the generated object should be a graph that meets certain constraints that are specified by a blackbox function such as in the design of photonic crystals (Christensen et al., 2020), crystal structure prediction (Kim et al., 2020), graph-constrained house generation (Tang et al., 2023), and general house plan generation (Nauata et al., 2021).
-In these settings, generating infeasible objects is penalized, and the generative model often has an inductive bias in the network architecture to generate feasible objects.
-These methods can use historical training datasets and handle unstructured objects; however, the output is often infeasible.
+Further work has modified deep generative models such as GAN (Goodfellow et al., 2014) to penalize constraint violation.
+These methods can generate objects respecting constraint graphs (Para et al., 2021), and blackbox constraints (Di Liello et al., 2020).
+Specialized approaches consider blackbox-constrained graphs such as in the design of photonic crystals (Christensen et al., 2020), crystal structure prediction (Kim et al., 2020), and house generation (Tang et al., 2023; Nauata et al., 2021).
+Here, feasibility is encouraged through penalties and inductive bias but is not guaranteed.
+Figure 1: GenCO diagram.
+A neural generator projects noise to a latent space which then gets used by a solver to generate provably feasible solutions.
+The solutions are then penalized with generative losses like Wasserstein distance (WGAN) or reconstruction (VQVAE) which can be backpropagated through the full pipeline.
+Additionally, GenCO can optimize individual objectives.
 Optimization-based priors
-In a related direction, previous work proposed conditioning variational autoencoders (VAE) (Kingma & Welling, 2013) with combinatorial programs (Misino et al., 2022).
-Here, the latent information is extracted by the encoder and then fed through a logic program, made differentiable via DeepProblog (Manhaeve et al., 2018), such as performing handwritten operations on digits.
+Previous work proposed conditioning VAE (Kingma & Welling, 2013) with combinatorial programs (Misino et al., 2022).
+Here, the latent information is extracted by the encoder and then fed through DeepProblog, a differentiable logic program (Manhaeve et al., 2018).
 The result is then fed through the decoder to generate the original image based on logical relationships.
-At test time, the goal is to generate objects that hopefully satisfy the logical relationship because the model was conditioned to do so.
-Here, the logical program helps to condition the generative model to generate objects based on patterns resulting from logical relationships and penalize cases where the logic is incorrect.
-Here, the generated objects are not guaranteed to satisfy the combinatorial constraints, but rather, the generative model has a logical structural prior.
-Previous work has made various optimization problems differentiable, such as quadratic programs (Amos & Kolter, 2017), probabilistic logic programs (Manhaeve et al., 2018), linear programs (Wilder et al., 2019a; Mandi & Guns, 2020; Elmachtoub & Grigas, 2017; Liu & Grigas, 2021), Stackelberg games (Perrault et al., 2020), normal form games (Ling et al., 2018), kmeans clustering (Wilder et al., 2019b), maximum likelihood computation (Niepert et al., 2021), graph matching (Rolínek et al., 2020), knapsack (Demirovic et al., 2019b; a), maxsat (Wang et al., 2019), mixed integer linear programs (Mandi et al., 2020; Paulus et al., 2021; Ferber et al., 2020), blackbox combinatorial solvers (Pogančić et al., 2020; Mandi et al., 2022; Berthet et al., 2020), nonlinear programs (Donti et al., 2017), continuous constraint satisfaction (Donti et al., 2020), cone programs (Agrawal et al., 2019).
-General-purpose methods are presented for minimizing downstream regret by learning surrogate loss functions (Shah et al., 2022; 2023; Zharmagambetov et al., 2023).
+At test time, the model generates objects that hopefully satisfy the logical relationship through model conditioning; however, there is no guarantee of constraint satisfaction.
+Previous work has made various optimization problems differentiable, such as quadratic programs (Amos & Kolter, 2017), probabilistic logic programs (Manhaeve et al., 2018), linear programs (Wilder et al., 2019a; Mandi & Guns, 2020; Elmachtoub & Grigas, 2017; Liu & Grigas, 2021), Stackelberg games (Perrault et al., 2020), normal form games (Ling et al., 2018), kmeans clustering (Wilder et al., 2019b), maximum likelihood computation (Niepert et al., 2021), graph matching (Rolínek et al., 2020), knapsack (Demirovic et al., 2019b, a), maxsat (Wang et al., 2019), mixed integer linear programs (Mandi et al., 2020; Paulus et al., 2021; Ferber et al., 2020), blackbox combinatorial solvers (Pogančić et al., 2020; Mandi et al., 2022; Berthet et al., 2020), nonlinear programs (Donti et al., 2017), continuous constraint satisfaction (Donti et al., 2020), cone programs (Agrawal et al., 2019).
+General-purpose methods are presented for minimizing downstream regret by learning surrogate loss functions (Shah et al., 2022, 2023; Zharmagambetov et al., 2023).
 Additionally, previous work has employed learnable linear solvers to solve nonlinear combinatorial problems (Ferber et al., 2023).
-Finally, a survey investigates the intersection between machine learning and optimization (Kotary et al., 2021).
+Finally, there is a survey on machine learning and optimization (Kotary et al., 2021).
 These approaches focus on identifying a single solution rather than generating diverse solutions.
-Additionally, many of these approaches are amenable for use by GenCO if a specific optimization problem better suits the generative problem.
+Additionally, many of these approaches can be integrated into GenCO if a specific optimization problem better suits the generative problem.
 Enhacing combinatorial optimization with generative models
 Recently, generative models have been proposed to improve combinatorial optimization.
-Zhang et al.
-(2022) use generative flow networks (gflownet) for robust scheduling problems.
-Sun & Yang (2023) use graph diffusion to solve combinatorial problems on graphs.
-Ozair et al.
-(2021) use vector quantized variational autoencoders to compress the latent space for solving planning problems in reinforcement learning.
-Zhao & You (2020) use generative adversarial networks to generate settings for sample average approximation in robust chance-constrained optimization.
+(Zhang et al., 2022) use generative flow networks (gflownet) for robust scheduling problems.
+(Sun & Yang, 2023) use graph diffusion to solve combinatorial problems on graphs.
+(Ozair et al., 2021) use vector quantized variational autoencoders to compress the latent space for solving planning problems in reinforcement learning.
+(Zhao & You, 2020) use generative adversarial networks to generate settings for sample average approximation in robust chance-constrained optimization.
 Additionally, in (Lopez-Piqueres et al., 2023), the authors generate continuous objects with tensor networks that satisfy linear constraints for optimization problems.
-These approaches target solving a fully specified optimization problem rather than generating objects with combinatorial constraints in mind.
+These approaches solve fully specified optimization problems rather than generating objects with combinatorial constraints.
 3 GenCO: Method Description
-Figure 1: An overview of GenCO.
-The top part of the diagram corresponds to the ordinary “GAN-style” generative model (chosen as an example).
-By integrating the “combinatorial loss” into both the loss function and the object generation process (bottom part), we effectively encode the combinatorial nature of the problem.
-3.1 Mathematical Formulation
-The key distinction in our framework centers around the incorporation of a combinatorial loss, a departure from conventional deep generative models like Generative Adversarial Networks (GANs) illustrated in Figure 1.
-Rather than directly generating the ultimate object of interest, be it images or any other complex entity, our generator $G{}(\epsilon{};\theta{})$ parameterized by $\theta{}$ takes random noise $\epsilon{}$ and transforms it into a problem representation $c$ that encapsulates the underlying problem’s essential features.
-This representation serves as a pivotal intermediary step.
-Subsequently, we feed $c$ to both the combinatorial loss and the generator loss (e.g.
-derived via a discriminator), ensuring that the solutions produced by the generator exhibit the desired characteristics: diversity, realism, and adherence to combinatorial properties.
-To train such a generator, we define the following bi-objective function, effectively balancing the generative loss $\mathcal{L}_{\text{gen}}{}$ with the combinatorial loss $\mathcal{C}{}$:
-$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(c)+\gamma\mathcal{C}{}(h(c))]$ (1)
-s.t. $c=G(\epsilon;\theta{}),$
-where $\mathcal{L}_{\text{gen}}{}$ encapsulates the generative loss, responsible for promoting diversity and realism.
-Its form can vary substantially depending on the specific generative model employed.
-For example, a Wasserstein GAN might utilize the adversarial discrepancy between outputs, while a Variational Autoencoder (VAE) would make use of the Evidence Lower Bound (ELBO).
-Conversely, the combinatorial loss $\mathcal{C}{}(\cdot)$, penalized by the weighting factor $\gamma$, is coupled with the downstream optimization solver $h(\cdot)$.
-This solver plays a crucial role in enforcing or encouraging combinatorial properties.
-For instance, one might seek to generate objects that must adhere to physical or logical constraints, all while minimizing a specified (linear) cost.
-In such a scenario, it becomes imperative to formulate an instance of a Mixed-Integer Linear Program (MILP) or another type of optimization problem with a well-defined feasible region.
-This problem instance then integrates seamlessly into the training pipeline, serving as the backbone of $\mathcal{C}{}$.
-Below, we consider two potential and practical settings for $\mathcal{C}{}$, each tailored to address specific real-world applications.
-3.2 Constrained Generator
-In this setting, we consider a problem domain where the goal is to train a generator that outputs feasible solutions, which natively handles diversity.
-To generate objects that are guaranteed to be feasible, we set $\gamma\to+\infty$ in equation 1 for any $c$ that does not satisfy constraints (defined by feasible region $\Omega$), 0 otherwise.
-To handle this directly with such a high value of $\gamma$ is computationally inefficient.
-Instead, we employ a ”projection layer” denoted as $h(\cdot)$ that maps $c$ to its nearest feasible point in the feasible set $\Omega$.
-Specifically, we define $h(c)=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$  [^1] Although a dot product is a natural choice of proximity for cosine distance, one could employ other metrics, e.g.
-$|c-x|$, where $c$ is an output of the generator $G(\epsilon;\theta)$.
-In this context, a generated problem description $c$ (as in Fig.1) can be viewed as an unconstrained generated object that is subsequently mapped via $h(\cdot)$ to $x$ to adhere to the defined constraints.
-With this approach, we can reformulate equation 1 into the following unconstrained optimization problem:
-$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(h(G(\epsilon;\theta{})))].$ (2)
-Note that the feasibility within region $\Omega$ is guaranteed to be satisfied due to the form of $h(\cdot)$.
-Directly minimizing this objective is possible thanks to the recent advancements in differentiable solvers literature (Agrawal et al., 2019; Pogančić et al., 2020; Sahoo et al., 2022).
-Algorithm 1 provides a high-level description of our method for this setting.
-It starts by initializing the generator parameters.
-In each iteration, a problem description is randomly selected using the generator.
-An optimization problem is then solved to find the solution that maximizes the similarity between $c$ within a defined feasible set ($x$ is the corrected/feasible version of $c$).
-Finally, the generator parameters are updated through backpropagation.
-We also provide an application of this generic algorithm to the concrete generative model examples of GANs and VQVAEs in Appendices B-C.
+Experiment Sol ${\bm{x}}$ Feas $\Omega$ Group Loss $\mathcal{L}$ Ind. Loss $\mathcal{D}$ Latent ${\bm{c}}$ Solver ${\bm{g}}$
+Game Design Game Level Playability WGAN (comb) Soft sol. Gurobi (ILP)
+Path Planning Min Path Path WGAN Min path (comb) RGB map Gurobi (LP)
+Photonic Device 0/1 Grid Manufacturing VQVAE (comb) Maxwell’s sim. (comb) Soft sol. Domain spec.
+Table 2: Summary of experimental settings highlighting the broad applicability of GenCO to settings with various optimizers, generative models, and loss functions.
+Losses operating on combinatorial objects rather than continuous latent space are highlighted with (comb).
+In the GenCO framework ( Figure 1), instead of direct optimization of the solution ${\bm{x}}$, we learn a distribution over a latent design space $\mathcal{C}$ so that for any ${\bm{c}}\in\mathcal{C}$, we get a solution ${\bm{x}}={\bm{g}}({\bm{c}})\in\mathcal{X}$ that is guaranteed to be feasible.
+Here ${\bm{g}}(\cdot)$ is a combinatorial solver that uses its input ${\bm{c}}$ to specify the combinatorial optimization problems that it solves.
+For example, ${\bm{g}}({\bm{c}}):=\operatorname*{arg\,min}_{{\bm{x}}\in\Omega}{\bm{c}}^{\top}{\bm{x}}$ outputs the solution to a mixed integer linear program (MILP) with coefficients ${\bm{c}}$.
+${\bm{g}}(\cdot)$ can also output a shortest path (via Dijkstra’s algorithm), given a weighted graph represented by ${\bm{c}}$, etc.
+During the training, we learn a generative model $G_{\theta}$ so that ${\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$ leads to solutions that minimize the loss:
+$\min_{\mathcal{C}:=\{{\bm{c}}_{j}\}}\mathcal{L}(\{{\bm{g}}({\bm{c}}_{j})\})+\gamma\sum_{j}\mathcal{D}({\bm{g}}({\bm{c}}_{j}))\quad\mathrm{s.t.}\ {\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$ (2)
+Here $\epsilon_{j}$ is random noise driving generation. This can be written as the following (with $\mathcal{X}:=\{{\bm{x}}_{j}\}$ representing a group of generated solutions):
+$\min_{\theta}\mathcal{L}(\mathcal{X})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})\quad\mathrm{s.t.}\ {\bm{x}}_{j}={\bm{g}}(G_{\theta}(\epsilon_{j}))$ (3)
+The generative model parameters $\theta$ can be learned via backpropagation through the combinatorial solver ${\bm{g}}(\cdot)$.
+Choices of group loss $\mathcal{L}$.
+The group loss $\mathcal{L}$ encourages the generated examples to capture the data distribution and can be instantiated using the Wasserstein distance as in Wasserstein GAN (Arjovsky et al., 2017), or Evidence Lower Bound (ELBO) as in VQVAE (Van Den Oord et al., 2017).
+In the WGAN case, the adversarial discriminator is trained only on examples satisfying the combinatorial constraints (${\bm{x}}$) rather than continuous solutions.
+Similarly, in VQVAE, the known feasible solutions are passed through the autoencoder followed by the solver ${\bm{g}}$ to ensure that the autoencoder loss is only computed on feasible solutions ${\bm{x}}$ rather than the continuous latent space ${\bm{c}}$.
+Choices of individual loss $\mathcal{D}$.
+We design the individual loss $\mathcal{D}$ to control the quality of individual solutions (e.g., the manufacture cost of a generated molecule ${\bm{x}}_{j}$ should be minimized).
+In many settings, this individual loss is nontrivial to encode in combinatorial solvers like Gurobi as in (Ferber et al., 2023).
+Loss on solution space $\Omega$ vs latent space $\mathcal{C}$ Note that to optimize GenCO’s objective ( Equation 3), we would need to run the combinatorial solver at every iteration, which involves calling MILP solvers and can be slow.
+On the other hand, if the group loss $\mathcal{L}$ or individual loss $\mathcal{D}$ can be evaluated in the latent space $\mathcal{C}$, then this extra cost is not needed during training.
+This occurs in the map generation experiments ( subsection 4.3), in which an observation ${\bm{c}}_{i}$ is provided (as an RGB image) and the optimal solution ${\bm{x}}_{i}$ can be computed accordingly (e.g., via an edge cost estimator and shortest path solver):
+$\min_{\theta}$ $\mathcal{L}(\mathcal{C})+\gamma\sum_{j}\mathcal{D}({\bm{x}}_{j})$ (4)
+$\ \ \mathrm{s.t.}$ $\ {\bm{x}}_{j}={\bm{g}}({\bm{c}}_{j})\quad\text{ and }\quad{\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$
+This formulation is related to SurCo-zero (Ferber et al., 2023), in which a latent representation ${\bm{c}}$ is optimized so that its inferred solution ${\bm{x}}={\bm{g}}({\bm{c}})$ by the solver ${\bm{g}}$ minimizes a nonlinear loss function $\mathcal{D}$.
+Mathematically, the training objective of SurCo is $\min_{\bm{c}}\mathcal{D}({\bm{g}}({\bm{c}}))$.
+When there are additional problem descriptions ${\bm{y}}_{i}$ (i.e., SurCo-prior), the latent design vector ${\bm{c}}_{i}={\bm{c}}_{\theta}({\bm{y}}_{i})$ becomes a function of ${\bm{y}}_{i}$ and thus the SurCo objective becomes $\min_{\theta}\sum_{i}\mathcal{D}({\bm{g}}({\bm{c}}_{\theta}({\bm{y}}_{i}));{\bm{y}}_{i})$.
+The main difference here is that GenCO considers the group loss $\mathcal{L}$ (e.g., matching the data distribution) on newly generated solutions, while SurCo focuses on the quality of solutions to known problem descriptions during training.
+In addition to  Figure 1, we summarize our proposed method in Algorithm 1.
+In each iteration, a set of solutions is generated using generator $G_{\theta}$ and a solver ${\bm{g}}$ (i.e., forward pass).
+We then evaluate the loss function using equation 3 (or equation 4).
+The parameters of $G_{\theta}$ are updated via backpropagation.
+Here, the differentiation through the combinatorial solver is achieved approximately by employing methods outlined in (Pogančić et al., 2020; Sahoo et al., 2022).
+Assumptions
+GenCO has several assumption on the individual loss $\mathcal{D}$, group loss $\mathcal{L}$, solver $g$, and latent space $c$.
+We assume the individual loss $\mathcal{D}$ is a function that maps a fixed solution to a loss value and provides gradient information useful for gradient descent.
+Note that if gradient information is not available, a neural network estimator may be learned to approximate the individual loss function.
+The group loss $\mathcal{L}$ imposes a loss on the distribution of instances generated.
+With deep generative models, this group loss is generally imposed by using learning-based proxies that act on individual examples.
+The assumption overall is that $\mathcal{L}$ can be optimized via stochastic gradient descent on the generated examples.
+We assume the solver $g$ is a function that maps from generated latent information $c$ to a feasible solution $x\in\Omega$.
+Additionally, we assume that previous work has made $g$ differentiable as a function of $c$ with various techniques proposed such as those described in related work.
+Finally, we assume the latent space $c$ simply represents values that can be mapped to via a neural network, and that the solver is well defined on that input.
   Output: Trained generator
-  Initialize generator parameters $\theta$;
+  Initialize generator parameters $\theta$
   while not converged do
-     Sample a noise $\epsilon$;
-     Sample a problem description $c\sim G(\epsilon;\theta)$;
-     Call a solver $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x=h(c)$;
-     Compute $\text{loss}=\mathcal{L}_{\text{gen}}{}(x^{*})$;
-     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$;
+     for $j=1\ldots K$ do
+        Sample a noise $\epsilon_{j}$
+        Sample a latent description ${\bm{c}}_{j}=G_{\theta}(\epsilon_{j})$
+        Call a solver ${\bm{x}}_{j}={\bm{g}}({\bm{c}}_{j})$
+     end for
+     Compute loss using  Equation 3 with $\mathcal{X}:=\{{\bm{x}}_{j}\}$ and $\mathcal{C}:=\{{\bm{c}}_{j}\}$.
+     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$
   end while
-Algorithm 1 Generator Training in Constrained Setting
-3.3 Penalized Generator
-The formulation from the previous section 3.2 naturally extends to penalty form, i.e., we encourage generated problem representation $c$ to respect certain combinatorial properties.
-In this scenario, we can directly leverage equation 1 as there is no need in projection layer:
-$\min_{\theta{}}\quad$ $\mathbb{E}_{\epsilon{}}[\mathcal{L}_{\text{gen}}{}(G(\epsilon;\theta{}))+\gamma\mathcal{C}{}(h(G(\epsilon;\theta{})))].$ (3)
-Here, we seamlessly integrate the generator into the objective function.
-While it is true that we do not impose any explicit constraints on the output of the generator, and one could use the output of $G$ directly, we contend that introducing a combinatorial penalty or loss as a penalty remains a crucial framework with significant practical utility.
-This approach serves to further guide the generation process and injects domain-specific knowledge that is easy to interpret, ensuring that the generated solutions exhibit desired combinatorial properties, a facet that holds immense practical relevance.
-For instance, we may aim to generate images tailored for path planning in strategy games.
-These maps feature various types of terrains, each incurring distinct costs when placed on the map.
-For efficient path planning, our goal is to find a “low cost” shortest path between two endpoints.
-Therefore, in this scenario, $\mathcal{C}$ incorporates a domain specific (fixed) mapping (e.g.
-a neural net) that transforms an image into a graph with corresponding weights assigned to each edge, followed by solving the shortest path problem (e.g.
-using Dijkstra’s algorithm).
-However, in this particular case, there is no need to introduce a “projection” operator as in the previous case since there are no explicit constraints that must be strictly satisfied.
-We can directly employ gradient-based optimization, allowing for differentiation through the combinatorial problem.
-This is achieved approximately by employing methods outlined in (Pogančić et al., 2020; Sahoo et al., 2022).
-Pseudocode 2 describes the application of GenCO for penalty setting.
-Furthermore, Appendix A.1 provides an application of this generic algorithm to a concrete example of GANs.
-  Output: Trained generator
-  Initialize generator parameters $\theta$;
-  (optional) Define a fixed mapping $g(\cdot)$ for $c$
-  while not converged do
-     Sample a noise $\epsilon$;
-     Sample a problem description $c\sim G(\epsilon;\theta)$;
-     (optional) apply a mapping $g(\cdot)$ to transform $c$;
-     Call a solver $x^{*}=h(g(c))$ ;
-     Compute $\text{loss}=\mathcal{L}_{\text{gen}}{}(c)+\gamma\mathcal{C}{}(x^{*})$;
-     Backpropagate $\nabla_{\theta}\text{loss}$ to update $\theta$;
-  end while
-Algorithm 2 Generator Training in Penalty Setting
-It is crucial to emphasize that the given penalized setting deviates from the conventional decision-focused learning (DFL) or smart predict-then-optimize paradigms (Elmachtoub & Grigas, 2017; Donti et al., 2017).
-In our approach, we do not possess access to the ground truth costs $c$, and the ultimate loss function encompasses both the generator’s loss and the output of a solver.
+Algorithm 1 GenCO
 4 Experiments
-We tested our approach on two applications: level generation for the Zelda game and map generation for path planning in Warcraft.
-Although both applications are in the game domain, they have quite different use cases.
-Nevertheless, both settings involve combinatorial optimization in the pipeline, which makes the typical use of deep generative models nontrivial.
-In these experiments, we demonstrated that GenCO significantly outperforms the baselines, efficiently finding diverse, realistic solutions that obey combinatorial properties.
-This success paves the way for combining combinatorial optimization with deep generative models.
-4.1 Generating game levels using GANs
-We evaluate our model on the task of generating diverse Zelda game levels. In this setting, we use GenCO to train using a GAN-like constrained formulation (refer to section 3.2 and Appendix B).
-4.1.1 Settings
-Explicitly Constrained GAN: Game Level Design
-We train GenCO on the task of generating Zelda game levels.
-Here, we are given examples of human-crafted game levels and are asked to generate fun new levels.
-The generated levels must be playable in that the player must be able to solve them by moving the character through a route that reaches the destination.
-Additionally, the levels should be realistic in that they should be somewhat similar to the real game levels as measured by an adversary that is trained to distinguish between real and fake images ($\mathcal{L}_{\text{gen}}{}$).
-We use the same dataset as (Zhang et al., 2020), consisting of 50 Zelda game levels, as well as the same network architectures, as we don’t tune the hyperparameters for our model in particular but rather compare the different approaches, all else equal.
-Here, we evaluate several approaches, including the previous work (Zhang et al., 2020), which we call GAN + MILP fix.
-This approach first trains a standard Wasserstein GAN architecture to generate the game levels.
-Specifically, they train the WGAN by alternatively training two components: a generator and a discriminator.
-The generator is updated to “fool” the discriminator as much as possible in that it tries to maximize the loss of the discriminator.
+We test GenCO on three applications: Zelda game level design, Warcraft path planning map generation, and diverse inverse photonic device design.
+Settings are summarized in Table 2, and all settings involve combinatorial optimization, making the application of deep generative models nontrivial.
+In these experiments, GenCO significantly outperforms the baselines, efficiently finding diverse, realistic solutions that obey combinatorial properties, paving the way for combining combinatorial optimization with deep generative models.
+4.1 Metrics
+We evaluate generative performance with density and coverage (Naeem et al., 2020).
+Density measures, for the average fake data point, how many real examples are nearby.
+Coverage measures how many real data points have a fake data point nearby.
+Formally, $M$ fake examples $Y_{j}$, are evaluated against $N$ real examples $X_{i}$.
+The metrics use the ball around $X_{i}$, with radius $r=\text{NND}_{k}(X_{i})$ being the maximimum distance from $X_{i}$ to its $k$ nearest real neighbors, to quantify generated density and coverage of the real data distribution.
+They are expressed as density = $\frac{1}{kM}\sum_{j=1}^{M}\sum_{i=1}^{N}1_{Y_{j}\in B(X_{i},\text{NND}_{k}(X_{i}))}$, and coverage = $\frac{1}{N}\sum_{i=1}^{N}1_{\exists j\text{s.t.}Y_{j}\in B(X_{i},\text{NND}_{k}(X_{i}))}$.
+4.2 Explicitly Constrained GAN: Game Level Design
+Approach % Unique $\uparrow$ Density $\uparrow$ Coverage $\uparrow$ GAN loss ($\mathcal{L}$) $\downarrow$ GenCO adversary $\downarrow$
+GAN + MILP fix (previous) 0.52 0.07 0.94 0.22 0.24
+GenCO - Fixed Adversary 0.22 0.05 0.98 -1.45 -0.85
+GenCO - Updated Adversary 0.995 0.06 0.82 -10.10 -4.49
+Table 3: Game level design comparison between the previous MILP postprocessing work, GenCO with an updating adversary, and GenCO with a fixed adversary, with all approaches guaranteed to generate feasible levels.
+We measure % Unique Density and Coverage to estimate diversity and distribution alignment.
+We also evaluate the adversary’s prediction to gauge how well a neural network can distinguish the generated levels from real levels.
+GAN+MILP Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+GenCO fixed Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+GenCO upd Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 2: Generated zelda level examples. “GenCO Updated” obtains solutions that seem more realistic than the empty MILP postprocessed GAN levels and GenCO Fixed Adversary levels.
+We evaluate GenCO on generating diverse and realistic Zelda game levels.
+Automatic level generation with playability guarantees helps designers create games with automatically generated content or iterate on development using a few examples.
+In this setting, we use GenCO to train using a constrained WGAN formulation.
+Here, we are given examples of human-crafted game levels $\mathcal{X}_{\mathrm{train}}$ and are asked to generate fun new levels.
+The generated levels must be playable in that the player must be able to complete them by moving the character through a route that reaches the destination $({\bm{x}}\in\Omega)$.
+Additionally, the levels should be realistic in that they should be similar to the real game levels as measured by an adversary that is trained to distinguish between real and fake images.
+This adversary imposes the generative group loss $\mathcal{L}{}$, effectively approximating the Wasserstein distance (Arjovsky et al., 2017).
+We use the same 50 Zelda levels as (Zhang et al., 2020), and the same neural network architecture, as we don’t tune the hyperparameters for our model in particular but rather compare the different approaches, all else equal.
+4.2.1 Baselines
+Here, we evaluate GenCO against previous work (Zhang et al., 2020)(GAN + MILP fix).
+This approach trains a standard Wasserstein GAN to approximate human game levels.
+Specifically, they train the WGAN by alternatively training a generator and a discriminator.
+The generator is updated to “fool” the discriminator (maximize the loss of the discriminator).
 The discriminator tries to correctly separate the generated and real game levels into their respective classes.
-When the practitioner then wants to generate a valid level, this approach generates a level with the generator and then fixes it using a MILP formulation that finds the nearest feasible game level where proximity is determined by cosine distance, which is equivalent to minimizing a dot product distance.
-We evaluate two variants of GenCO, GenCO - Fixed Adversary, which iteratively updates the generative model to fool a fixed pretrained adversary, and GenCO - Updated Adversary, which updates both the generator and adversary during training.
-The fixed adversary approach trains to fool a pretrained adversary that is obtained from the fully trained GAN from previous work (Zhang et al., 2020).
-Both GenCO approaches are initialized with the fully trained GAN from previous work (Zhang et al., 2020).
+When the practitioner wants to generate a valid level, this approach generates a soft level and fixes it using a MILP that finds the nearest feasible game level in terms of cosine distance (dot product).
+We evaluate two variants of GenCO, GenCO - Fixed Adversary, which trains against a fixed pretrained adversary from previous work (Zhang et al., 2020), and GenCO - Updated Adversary, which updates both the generator and adversary during training.
+Both approaches are initialized with the fully trained GAN from previous work (Zhang et al., 2020).
+Importantly, the adversaries act on the generated solution ${\bm{x}}\in\Omega$, rather than the latent soft solution ${\bm{c}}$.
 4.2.2 Results
-Approach % unique
-GAN loss ($\mathcal{L}_{\text{gen}}{}$)
-(lower better)
-GenCO adversary
-(lower better)
-GenCO - Updated Adversary 0.995 -10.10 -4.49
-GenCO - Fixed Adversary 0.22 -1.45 -0.85
-GAN + MILP fix (Previous Work) 0.52 0.22 0.24
-Table 1: Game level design comparison.
-We compare GenCO with an updating adversary against GenCO with a fixed adversary, and lastly, previous work that postprocesses solutions to be valid game levels.
-Here, all levels are valid, but GenCO exhibits more diversity and realism.
-GAN + MILP Fix Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-GenCO Fixed Adv Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-GenCO Updated Adv Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 2: Generated zelda game examples. The GenCO Updated Adversary obtains solutions that seem more realistic than mainly empty fixed GAN instances and fixed adversary instances.
-We present a table of results in Table 1 and examples in Figure 2, which includes the performance of the previous approach as well as two variants of our proposed approach.
+Results are presented in Table 3 and examples in Figure 2, which includes the performance of the previous approach as well as two variants of GenCO.
 In these settings, we estimate performance based on sampling 1000 levels.
@@
 Furthermore, given that the levels are trained on only 50 examples, we can obtain many more game levels.
+Overall, GenCO - Updated Adversary obtains the best uniqueness and generative loss.
+However, it obtains neither the best density nor coverage since game levels are generally sparse (the majority of tiles are empty floors), meaning that having a slightly shifted map will result in a higher distance than a completely empty map.
+As a result, coverage and density have counterintuitive behaviors in discrete settings while in standard image domains, samples are continuous and unconstrained.
+Additionally, density and coverage are relatively low.
 Note here that both approaches are guaranteed to give playable levels as they are postprocessed to be valid.
-However, GenCO is able to generate more diverse solutions that are also of higher quality.
+However, GenCO is able to generate more diverse solutions that are also of higher realism quality.
+Approach Density $\uparrow$ Coverage $\uparrow$ GAN loss ($\mathcal{L}$) $\downarrow$ SP loss ($\mathcal{D}$) $\downarrow$
+Ordinary GAN 0.81 0.98 0.6147 36.45
+GAN + semantic loss (Di Liello et al., 2020) 1.09 0.98 0.8994 35.61
+GenCO (ours) 0.94 0.93 0.6360 23.99
+Table 4: Map generation for Warcraft performance comparison on 100 instances.
+The goal is to generate maps which are similar to the data distribution (Density / Coverage), realistic (lower GAN loss or group loss $\mathcal{L}$), and have small shortest paths (SP or individual loss $\mathcal{D}$).
 Uniqueness
-As shown in Table 1, GenCO, with an updated adversary, obtains the highest percentage of unique solutions, generating 995 unique solutions out of 1000.
-This is significantly higher than the previous work, which only generated 520 unique solutions.
+As shown in Table 3, GenCO, with an updated adversary, obtains the highest percentage of unique solutions, generating 995 unique solutions out of 1000, significantly higher than the 520 unique solutions in previous work.
 This is likely due to the fact that the generator is trained with the downstream fixing explicitly in the loop.
@@
 In this experiment, we consider somewhat different and more challenging for GenCO setting of generation of image maps for strategy games like Warcraft.
-This poses a greater challenge because the generator $G$ described in equation 1 is not restricted and can produce any image of a map (setting descrbied in section 3.3).
-However, we are required to encourage $G$ to generate a map with two main criteria: 1) it should resemble real and diverse game maps, as determined by an adversary trained to differentiate between real and fake images; 2) the cost of the shortest path from the top-left to the bottom-right corners (source and destination) should be minimized.
-Intuitively, it corresponds to (mostly) a diagonal part of the image, and a map can include various elements (terrains) like mountains, lakes, forests, and land, each with a specific cost (for example, mountains may have a cost of 3, while land has a cost of 0).
-To calculate this, we pass the generated image to the fixed ResNet (called “cost NN”) to get the graph representation together with edge costs.
-The cost NN corresponds to the mapping $g(\cdot)$ in Algorithm 2.
-This is then followed by the shortest path solver ($h(\cdot)$ in Algorithm 2).
-The objective is to populate the map with a diverse range of objects while ensuring the shortest path remains low-cost.
+This poses a greater challenge because the generator $G$ is not restricted and can produce any RGB image of a map (group loss $\mathcal{L}$ operates directly on latent space $\mathcal{C}$, i.e., $\mathcal{L}(\mathcal{C})$ in eq.
+3).
+However, we would like to encourage $G$ to generate a map with two main criteria: 1) it should resemble real and diverse game maps, as determined by the group Loss $\mathcal{L}(\mathcal{C})$; 2) the cost of the shortest path (SP) from the top-left to the bottom-right corners (source and destination) should be minimized (group loss $\mathcal{D}({\bm{g}}({\bm{c}}))$).
+Intuitively, the SP corresponds to (mostly) a diagonal part of the image, and a map can include various elements (terrains) like mountains, lakes, forests, and land, each with a specific cost (for example, mountains may have a cost of 3, while land has a cost of 0).
+To calculate this, we pass the generated image to the fixed ResNet to get the graph representation together with edge costs.
+This is then followed by the shortest path solver (${\bm{g}}$).
+The objective is to populate the map with a diverse range of objects while ensuring the shortest path remains low-cost: $\mathcal{L}(\mathcal{C})+\gamma\sum_{j}\mathcal{D}({\bm{g}}({\bm{c}}_{j}))$.
 We use the same dataset as in (Pogančić et al., 2020) with DCGAN architecture adapted from (Zhang et al., 2020).
+Note that the maps are unconstrained, representing the generator’s latent space.
+However, the generated feasible solution is a valid path through the map.
+Additionally, the generative loss acts on the latent space $c$ rather than the feasible solution as we are given maps rather than solutions.
+The resnet architecture is part of the problem formulation in that it determines costs.
 More implementation details and experimental settings can be found in Appendix A.
 4.3.1 Baselines
-We examine various baseline models, including an “Ordinary GAN” that does not incorporate the Shortest Path objective (see the top part of Fig.
-1).
-Like the Zelda experiment, this approach employs a standard Wasserstein GAN architecture closely following (Zhang et al., 2020).
+We examine various baseline models, including an “Ordinary GAN” that does not incorporate the Shortest Path objective.
+As in the Zelda experiment, this approach employs a standard Wasserstein GAN architecture closely following (Zhang et al., 2020).
 However, in this case, we generate images directly instead of encoding game levels.
@@
 The discriminator, on the other hand, endeavors to accurately distinguish between generated and authentic game levels and categorize them accordingly.
-The next baseline is “GAN + cost NN”, which simultaneously feeds the generator’s output into both the discriminator and a ResNet that calculates the costs for each edge in the grid.
-We then average the output to obtain the final loss for generator.
-We then average these outputs to derive the ultimate loss for the generator.
-While this approach considers costs associated with objects, it does not incorporate information about the Shortest Path.
+The next baseline is a “soft” penalty approach: “GAN + semantic loss”.
+It uses a fixed NN to extract representation of the image in the form of cost vector (obtained from the fixed ResNet).
+This vector is then averaged and added to the final objective as a regularization.
+The concept is similar to the semantic loss appeared in (Di Liello et al., 2020).
+The difference is that instead of building a circuit (via knowledge compilation) to encode the constraints, we use a fixed pre-trained ResNet as a surrogate to obtain the cost vector.
+While this baseline considers costs associated with objects, it does not incorporate information about the Shortest Path.
 As for the GenCO, it generalizes both of these approaches incorporating combinatorial solver into the pipeline.
@@
 4.3.2 Results
-Quantitative results are showcased in Table 2.
-The “Ordinary GAN” focuses exclusively on the GAN’s objective, without considering the Shortest Path’s objective $f$.
+Quantitative results are showcased in the Table 4.
+The “Ordinary GAN” focuses exclusively on the GAN’s objective (group loss $\mathcal{L}$), without considering the Shortest Path’s objective (individual loss $\mathcal{D}$).
 In contrast, GenCO achieves higher performance with regards to the SP’s objective, albeit with a slight reduction in GAN’s loss.
-On the other hand, the “GAN + cost NN” approach optimizes using the cost vector of the entire map (generated by ResNet).
-In other words, it tries to uniformly avoid placing costly objects in any part of the image, whereas we are interested only in the shortest path.
-While it demonstrates a modest enhancement in the objective $f$, it experiences a notable decline in terms of generator loss ($\mathcal{L}_{\text{gen}}$).
+On the other hand, the “GAN + semantic loss” approach tries to uniformly avoid placing costly objects in any part of the image, whereas we are interested only in the shortest path.
+While it demonstrates a modest enhancement in the objective $\mathcal{D}$, it experiences a notable decline in terms of group loss ($\mathcal{L}$).
 This indicates a trade-off between optimizing for the Shortest Path and the GAN’s objective.
-Approach
-GAN loss ($\mathcal{L}_{\text{gen}}$)
-(lower better)
-SP loss ($\mathcal{C}$)
-(lower better)
-Ordinary GAN 0.6147 36.45
-GAN + cost NN 0.8994 35.61
-GenCO (ours) 0.6360 23.99
-Table 2: Performance comparison on Warcraft map generation.
-Our here goal is to create a map that is both realistic (“fools” discriminator, i.e.
-lower GAN loss) and has small objective for the shortest path (SP).
-Results are averaged across 100 instances.
+Here, the density and coverage are comparable between all approaches.
+While GenCO doesn’t obtain the highest density and coverage, it does obtain good SP loss and comparable GAN loss with Ordinary GAN, indicating a trade-off between fitting the distribution and obtaining high-quality solutions.
 Such quantitative results directly translate into image qualities.
@@
 This makes sense since WGAN is trained with no information about grid costs.
-In contrast, the “GAN + cost NN” approach produces less costly maps, albeit with reduced diversity.
+In contrast, the “GAN + semantic loss” approach produces less costly maps, albeit with reduced diversity.
 For instance, most part of the image is populated with the same object.
@@
 The Shortest Path is efficient, and the map exhibits a rich variety of features simultaneously.
-Ordinary GAN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-GAN + cost NN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-GenCO (ours) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 3: A subset of generated Warcraft map images using various methods.
+$\gamma$ SP Loss ($\mathcal{D}$) $\downarrow$ Density $\uparrow$ Coverage $\uparrow$
+0 36.45 0.81 0.98
+1e-4 27.66 0.93 0.93
+1e-3 23.99 0.94 0.93
+3e-3 23.76 0.75 0.86
+5e-3 18.02 0.49 0.61
+1e-2 0.00 0.00 0.00
+Table 5: $\gamma$ ablation, trading off individual and generative loss. High trade-off parameters improve optimization performance while low trade-off values give slightly better generation.
+Additionally, in Table 5 we evaluate sensitivity to generative and shortest path loss trade-off parameter.
+While the shortest path loss improves with high trade-off parameter, the image quality metrics start to deteriorate rapidly.
+Eventually, our approach generates blank or empty images, as such images result in the shortest path with zero cost.
+This behavior could be attributed to the combinatorial nature of the problem.
+Ord. GAN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+GAN + sem. loss Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+GenCO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 3: A subset of generated Warcraft map images.
 “Ordinary GAN” generates “very costly” maps (e.g.
-mountains, lakes) along the Shortest Path from top-left to bottom-right; “GAN + cost NN” generates less costly maps but they are less diverse; GenCO: SP path is cheap and the map is diverse at the same time.
+mountains, lakes) along the Shortest Path from top-left to bottom-right; “GAN + semantic loss” generates less costly maps but they are less diverse; GenCO: SP path is cheap and the map is diverse at the same time.
+4.4 Constrained VAE: Photonic Design
+Approach % Unique $\uparrow$ Density $\uparrow$ Coverage $\uparrow$ Avg Solution Loss $\downarrow$
+VQVAE + postprocess 30.6% 0.009 0.006 1.244
+GenCO (reconstruction only) 100% 0.148 0.693 1.155
+GenCO (objective only) 46.6% 0.013 0.036 0
+GenCO (reconstruction + objective) 100% 0.153 0.738 0
+Table 6: Inverse photonic design comparison evaluating variants of GenCO using VQVAE against the same model architecture which postprocesses solutions.
+Solution loss is evaluated using a simulation of Maxwell’s equations and is optimal at 0.
+We finally evaluate GenCO for training a Vector Quantized Variational Autoencoder (VQVAE) to ensure that all decoded examples satisfy combinatorial manufacturing constraints ${\bm{x}}\in\Omega$, are each minimizers of a photonic loss $\mathcal{D}$, and are similar to given solutions via the ELBO loss $\mathcal{L}$.
+We evaluate GenCO against a generative + postprocess baseline (Zhang et al., 2020) on the inverse photonic design setting.
+The inverse photonic design problem (Schubert et al., 2022) asks how to design a device consisting of fixed and void space to route wavelengths of light from an incoming location to desired output locations at high intensity.
+Here the feasible region consists of satisfying manufacturing constraints that a die with a specific shape must be able to fit in every fixed and void space.
+Specifically, the fixed and void regions must represent the union of the die shape at different locations.
+The objective function here consists of a nonlinear but differentiable simulation of the light using Maxwell’s equations.
+Previous work demonstrated an approach for finding a single optimal solution to the problem.
+However, we propose generating a diverse collection of high-quality solutions using a dataset of known solutions.
+In practice, there are various applications for generating multiple feasible photonic devices.
+Firstly, there are many devices that globally minimize the objective function by matching the specification perfectly.
+With a generator, practitioners can select from a diverse array of solutions that already optimize both specification and satisfy manufacturing constraints, quickly generating more examples if needed.
+Upon inspecting the solutions, they may select one that satisfies their needs, or derive high-level insights from the diversity of viable solutions to improve the design process.
+Lastly, optimizing the encoded objective function over the feasible region may not necessarily result in the best solution considering unquantified components like aesthetics, further manual scrutiny, and simply tacit domain expertise.
+Regardless, the selected design must always be optimal and satisfy the problem constraints.
+As such, we seek to generate diverse optimal solutions to aid the practitioner when a single solution fails to capture the full range of possibilities.
+In this setting, we train a vector quantized variational autoencoder (VQVAE) (Van Den Oord et al., 2017).
+During training, the encoder encodes a known solution and then decodes an object with the same shape as the input space, using neural networks for the encoder and decoder.
+The continuous decoded object is then fed into the constrained optimization layer to enforce that the generated solution is feasible.
+This feasible solution is then used to calculate the reconstruction loss.
+Furthermore, in this setting we have an optimization-based objective function that consists of the simulation of Maxwell’s equations.
+As such, we ablate the data distribution approximation, and penalty components of GenCO: whether or not to train using the reconstruction loss, and whether or not to penalize generated solutions based on Maxwell’s equations.
+We consider a baseline of training the same generative architecture without combinatorial optimization and then postprocessing generated examples during evaluation using a combinatorial solver.
+At test time, we sample noise and decode to the original input space.
+We then obtain a feasible solution using the combinatorial solver.
+The results in Table 6 include the solution uniqueness rate, the average loss evaluated using Maxwell’s equations, as well as the density and coverage.
+The dataset is obtained by expensively running previous work (Ferber et al., 2023) until we obtain 100 optimal solutions.
+We evaluate performance on generating 1000 feasible examples.
+These results in Table 6 demonstrate that postprocessing obtains very few unique solutions which all have high loss and furthermore don’t resemble the data distribution.
+This is largely due to the method not being trained with the postprocessing end-to-end.
+Although it closely approximates the data distribution with continuous and infeasible objects, when these continuous objects are postprocessed to be feasible, they are no longer representative of the data distribution and many continuous solutions collapse to the same discrete solution.
+Additionally, GenCO obtains high density and coverage compared to the baseline postprocessed VQVAE indicating that training end-to-end helps in this setting.
+GenCO - reconstruction only gives many unique solutions that resemble the data distribution.
+However, the generated devices fail to perform optimally in the photonic task at hand.
+Disregarding the reconstruction loss and only training the decoder to generate high-quality solutions yields high-quality but redundant solutions.
+Combining both the generative reconstruction penalty as well as the nonlinear objective, GenCO generates a variety of unique solutions that optimally solve the inverse photonic design problem and are representative of the data distribution.
 5 Conclusion
@@
 The underlying idea of our framework is to combine the flexibility of deep generative models with the guarantees of optimization solvers.
-By training the generator end-to-end with a surrogate linear combinatorial solver, GenCO generates diverse and combinatorially feasible solutions, with the generative loss being computed only on feasible solutions.
-We have tested GenCO on various combinatorial optimization problems and generative settings, including GAN in Zelda game level generation and Warcraft map generation for path planning, demonstrating the flexibility of our framework for integrating into different generative paradigms.
-Our framework consistently produced diverse and high-quality solutions that satisfy the combinatorial constraints, which can be flexibly encoded using general-purpose combinatorial solvers.
-Appendix A Detailed experimental setup for map generation
+By training the generator end-to-end with a combinatorial solver, GenCO generates diverse and combinatorially feasible solutions, with the generative loss being computed only on feasible solutions.
+Currently, many of GenCO’s requirements are tied to those of the space of differentiable solvers which is a rapidly evolving field.
+Some of the key bottlenecks researchers are working to address are scalability, differentiability, and the range of optimization problems that can be differentiated.
+Another frontier of differentiable solver research is in learning constraints from data.
+Currently it is assumed that practitioners have access to a mathematical representation of the constraints that can be optimized over.
+However, in some domains such as molecular design, the constraints such as binding affinity are not available analytically and it is nontrivial to guarantee feasibility using mathematical programs.
+Learning mathematical representations of these constraints from data is an active area of research.
+Additionally, as with many generative models, it is important to keep real-world deployment in mind.
+Even though GenCO can guarantee feasibility with respect to explicitly encoded constraints, it cannot yet guarantee feasibility with respect to constraints that are desirable but not explicitly encoded.
+For instance, unless explicitly encoded or leveraging other alignment research, GenCO cannot guarantee fairness, privacy, or positive social impact.
+Similarly, in image generation, it may be difficult to encode some constraints such as ensuring that pixels representing hands have exactly 5 fingers using a diffusion model in image generation.
+Overall, GenCO is a step towards enforcing desirable constraints on the outputs of generative models, but it does not immediately solve all problems in ensuring the safety of generative models.
+Overall, we propose GenCO, the first generative model that generates diverse solutions, optimizing differentiable nonlinear objectives with guaranteed feasibility for explicitly combinatorial constraints.
+We test GenCO on three different scenarios: inverse photonic design, game level design, and path planning, showing that GenCO outperforms existing methods along multiple axes.
+We have tested GenCO on various combinatorial optimization problems and generative settings, including GAN in Zelda game level generation, Warcraft map generation for path planning, and inverse photonic design.
+These settings require that GenCO handle a variety of generative architectures (GAN / VAE), optimization problems (linear programs, quadratic programs, mixed integer programs), and solvers (SCIP, Gurobi, shortest path solvers, blackbox domain specific solvers), demonstrating the flexibility of our framework for integrating different generative paradigms and optimization methods.
+Our framework consistently produces diverse and high-quality solutions that satisfy combinatorial constraints, which can be flexibly encoded using general-purpose or domain specific combinatorial solvers.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning, deep generative models, and combinatorial optimization.
+There are many potential societal consequences of the fields, none which we feel must be specifically highlighted here.
+Additionally, while our work focuses on guaranteeing the output of generative models satisfies specified constraints, it may be difficult to characterize all types of constraints which may be impossible to specify or enumerate.
+Acknowledgment
+The research at the University of Southern California was supported by the National Science Foundation (NSF) under grant number 2112533. We also thank the anonymous reviewers for helpful feedback.
+Appendix A Detailed experimental setup for map generation (section 4.3)
 A.1 Pseudocode
-1:  Initialize generator parameters $\theta_{gen}$;
-2:  Initialize adversary parameters $\theta_{adv}$;
-3:  Input: distribution of the true dataset ($c_{true}\sim p_{\text{data}(c)}$), GAN’s objective $\mathcal{L}(\theta_{gen},\theta_{adv})$;
+1:  Initialize generator parameters $\theta_{gen}$
+2:  Initialize adversary parameters $\theta_{adv}$
+3:  Input: distribution of the true dataset (${\bm{c}}_{true}\sim p_{\text{data}({\bm{c}})}$), GAN’s objective $\mathcal{L}(\theta_{gen},\theta_{adv})$
 4:  for epoch $e$ do
 5:     for batch $b$ do
-6:        Sample a noise $\epsilon$;
-7:        Sample true examples from dataset $c_{true}\sim p_{\text{data}}(c)$;
-8:        Sample fake examples using $c_{fake}\sim G(\epsilon;\theta_{gen})$;
-9:        Transform $c_{fake}$ into the coefficients of the optimization problem: $c=g(c_{fake})$;
-10:        Solve: $x^{*}=\operatorname*{arg\,min}_{x\in\Omega}c^{T}z$;
-11:        Backpropagate $\nabla_{\theta_{gen}}\left[\mathcal{L}(c_{fake};\theta_{gen},\theta_{adv})+\beta c^{T}x^{*}\right]$ to update $\theta_{gen}$;
-12:        Backpropagate $\nabla_{\theta_{adv}}\left[-\mathcal{L}(c_{fake};\theta_{gen},\theta_{adv})+\mathcal{L}(c_{real};\theta_{gen},\theta_{adv})\right]$ to update $\theta_{adv}$;
+6:        Sample a noise $\epsilon$
+7:        Sample true examples from dataset $c_{true}\sim p_{\text{data}}(c)$
+8:        Sample fake examples using ${\bm{c}}_{fake}\sim G(\epsilon;\theta_{gen})$
+9:        Transform ${\bm{c}}_{fake}$ into the coefficients of the optimization problem: ${\bm{c}}=f({\bm{c}}_{fake})$
+10:        Solve: $x^{*}={\bm{g}}({\bm{c}})=\operatorname*{arg\,min}_{x\in\Omega}{\bm{c}}^{T}x$
+11:        Backpropagate $\nabla_{\theta_{gen}}\left[\mathcal{L}({\bm{c}}_{fake};\theta_{gen},\theta_{adv})+\beta c^{T}x^{*}\right]$ to update $\theta_{gen}$
+12:        Backpropagate $\nabla_{\theta_{adv}}\left[-\mathcal{L}({\bm{c}}_{fake};\theta_{gen},\theta_{adv})+\mathcal{L}({\bm{c}}_{real};\theta_{gen},\theta_{adv})\right]$ to update $\theta_{adv}$
 13:     end for
@@
 Subsequently, the synthetic data undergoes a fixed mapping (e.g.
-ResNet in our experiments), called cost neural net (or cost NN), to obtain coefficients for the optimization problem, specifically the edge weights for the Shortest Path.
+ResNet in our experiments), called cost neural net (or cost NN), to obtain coefficients for the optimization problem (the edge weights for the Shortest Path).
 Following this, we invoke a solver that provides us with a solution and its associated objective (lines 10–11).
-We then proceed to update the parameters of the generator $G$ using both the GAN’s objective and the solver’s objective.
+We then proceed to update the parameters of the generator $G$ using both the GAN’s objective (group loss $\mathcal{L}$) and the solver’s objective (individual loss $\mathcal{D}$).
 Finally, we refine the parameters of the adversary (discriminator) in accordance with the standard GAN’s objective.
 A.2 Settings
-We employ ResNet as the mapping $g(\cdot)$ from Algorithm 3, which transforms an image of a map into a $12\times 12$ grid representation of a weighted directed graph: $g:\Re^{96\times 96\times 3}\rightarrow\Re^{12\times 12}$.
-The first five layers of ResNet18 are pre-trained (75 epochs, Adam optimizer with lr=$5e-4$) using the dataset from Pogančić et al.
-(2020), comprising 10,000 labeled pairs of image–grid (refer to the dataset description below).
+We employ ResNet as the mapping $f(\cdot)$ from the above Algorithm 2, which transforms an image of a map into a $12\times 12$ grid representation of a weighted directed graph: $f:\Re^{96\times 96\times 3}\rightarrow\Re^{12\times 12}$.
+The first five layers of ResNet18 are pre-trained (75 epochs, Adam optimizer with lr=$5e-4$) using the dataset from (Pogančić et al., 2020), comprising 10,000 labeled pairs of image–grid (refer to the dataset description below).
 Following pretraining, we feed the output into the Shortest Path solver, using the top-left point as the source and the bottom-right point as the destination.
-The resulting objective value from the Shortest Path corresponds to $f$.
+The resulting objective value from the Shortest Path corresponds to ${\bm{g}}$.
 Dataset
-The dataset used for training in the Shortest Path problem with $k=12$ comprises 10,000 randomly generated terrain maps from the Warcraft II tileset Pogančić et al.
-(2020) (adapted from Guyomarch (2017)).
+The dataset used for training in the Shortest Path problem with $k=12$ comprises 10,000 randomly generated terrain maps from the Warcraft II tileset (Pogančić et al., 2020) (adapted from (Guyomarch, 2017)).
 These maps are represented on a $12\times 12$ grid, with each vertex denoting a terrain type and its associated fixed cost.
@@
 A.3 Architecture
-We employ similar DCGAN architecture taken from Zhang et al.
-(2020) (see fig.
+We employ similar DCGAN architecture taken from (Zhang et al., 2020) (see fig.
 3 therein).
@@
 The discriminator mirrors this architecture in reverse order.
-The entire structure is trained using the WGAN algorithm, as described inZhang et al.
-(2020).
+The entire structure is trained using the WGAN algorithm, as described in(Zhang et al., 2020).
 Appendix B GANs with combinatorial constraints
@@
 Furthermore, we can ensure that the objective function is optimized by penalizing the generator based on the generated solutions’ objective values:
-$\mathcal{L}_{\text{gen}}{}(G(\epsilon;\theta_{gen}))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(G(\epsilon;\theta_{gen})))\right]$ (4)
+$\mathcal{L}{}(G(\epsilon;\theta_{gen}))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(G(\epsilon;\theta_{gen})))\right]$ (5)
 where $f_{\theta_{adv}}$ is an adversary (a.k.a. discriminator) and putting this in the context of equation 1 leads to:
-$\min_{\theta_{gen}}\quad$ $\mathcal{L}_{\text{gen}}{}(S(\tilde{G}(\epsilon;\theta_{gen})))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(S(\tilde{G}{}(\epsilon;\theta_{gen}))))\right]$ (5)
+$\min_{\theta_{gen}}$ $\mathcal{L}{}(S(\tilde{G}(\epsilon;\theta_{gen})))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(S(\tilde{G}{}(\epsilon;\theta_{gen}))))\right]$ (6)
 where $\tilde{G}{}$ is unconstrained generator and $S$ is a surrogate combinatorial solver as described above.
@@
 However, that part does not depend on combinatorial solver and can be trained as in usual GAN’s.
-The algorithm is presented in pseudocode 4.
-  Initialize generator parameters $\theta_{gen}$;
-  Initialize adversary parameters $\theta_{adv}$;
+The algorithm is presented in pseudocode 3.
+  Initialize generator parameters $\theta_{gen}$
+  Initialize adversary parameters $\theta_{adv}$
   for epoch $e$ do
      for batch $b$ do
-        Sample problem $\epsilon{}$;
-        Sample true examples from dataset $x_{true}\sim p_{\text{data}}(x)$;
-        Sample linear coefficients $c\sim G(\epsilon{};\theta_{gen})$;
-        Solve $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$;
-        Backpropagate $\nabla_{\theta_{gen}}\mathcal{L}_{\text{gen}}{}(x^{*};\theta_{gen})$ to update generator (equation 5);
-        Backpropagate $\nabla_{\theta_{adv}}\left[\log(f_{\theta_{adv}}(x_{true}))-\mathcal{L}_{\text{gen}}{}(x^{*};\theta_{gen})\right]$ to update adversary;
+        Sample problem $\epsilon{}$
+        Sample true examples from dataset $x_{true}\sim p_{\text{data}}(x)$
+        Sample linear coefficients $c\sim G(\epsilon{};\theta_{gen})$
+        Solve $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$
+        Backpropagate $\nabla_{\theta_{gen}}\mathcal{L}{}(x^{*};\theta_{gen})$ to update generator (equation 6)
+        Backpropagate $\nabla_{\theta_{adv}}\left[\log(f_{\theta_{adv}}(x_{true}))-\mathcal{L}{}(x^{*};\theta_{gen})\right]$ to update adversary
      end for
   end for
-Algorithm 4 GenCO GenCO in the constrained generator setting
+Algorithm 3 GenCO in the constrained generator setting
 Appendix C GenCO – VQVAE
@@
 Here, we simply train VQVAE on a dataset of known objective coefficients, which solves the problem at hand.
-A variant of this also puts the decision-focused loss on the generated objective coefficients, running optimizer $g_{\text{solver}}$ on the objective coefficients to get a solution and then computing the objective value of the solution.
-$L_{\text{ELBO}}(c,\theta,E)=\mathbb{E}_{q_{\theta}(z|c)}[\log p_{\theta}(c|z)]-\beta\cdot D_{\text{KL}}(q_{\theta}(z|c)||p(z))+\gamma\cdot\|sg(\mathbf{e}_{k})-\mathbf{z}_{e,\theta}\|_{2}^{2}$ (6)
-Here $z$ is an embedding vector, $c$ is the objective coefficients, $\log p_{\theta}(c|z)$ is a loss calculated via the mean squared error between the decoder output and the original input objective coefficients, $q_{\theta}(z|c)$ is the encoder, $p(z)$ is the prior, and $sg(\cdot)$ is the stop gradient operator, $E$ is a discrete codebook that is used to quantize the embedding.
-$\mathcal{L}_{\text{optimization}}=\mathbb{E}_{c\sim p_{\theta}(c\mid z)}\left[f_{\text{obj}}(g_{\text{solver}}(c;y))\right]$ (7)
+A variant of this also puts the decision-focused loss on the generated objective coefficients, running optimizer $g$ on the objective coefficients to get a solution and then computing the objective value of the solution.
+$\mathcal{L}(\mathcal{X})=\text{ELBO}({\bm{x}},\theta,E)=\mathbb{E}_{q_{\theta}(z|{\bm{x}})}[\log p_{\theta}({\bm{x}}|z)]-\beta\cdot D_{\text{KL}}(q_{\theta}(z|{\bm{x}})||p(z))+\gamma\cdot\|sg(\mathbf{e}_{k})-\mathbf{z}_{e,\theta}\|_{2}^{2}$ (7)
+Here $z$ is an embedding vector, $c$ is the objective coefficients, $\log p_{\theta}({\bm{x}}|z)$ is a loss calculated via the mean squared error between the decoder output and the original input objective coefficients, $q_{\theta}(z|{\bm{x}})$ is the encoder, $p(z)$ is the prior, and $sg(\cdot)$ is the stop gradient operator, $E$ is a discrete codebook that is used to quantize the embedding.
+$\mathcal{D}=\mathbb{E}_{c\sim p_{\theta}(c\mid z)}\left[f_{\text{obj}}(g(c;y))\right]$ (8)
 The algorithm below maximizes a combination of the losses in Equation equation 7 and Equation equation 8.
-Input:Training data distribution $D$ over problem info $y$ and known high-quality objective coefficients $c$, regularization weight $\beta$, linear surrogate solver $g_{\text{solver}}$, nonlinear objective $f_{\text{objective}}$.
-Output:Trained encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and codebook $E$ Initialize the parameters of the encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and the codebook $E=\{e_{1},e_{2},\ldots,e_{K}\}$ with $K$ embedding vectors;
-for $t=1$ to $T$ do
-       Sample $y,c$ from the distribution $D$;
-       Compute the encoder output $z_{e}=f_{\text{enc}}(y,c)$;
-       Find the nearest embedding vector $z_{q}=\arg\min_{e\in E}\|z_{e}-e\|_{2}^{2}$;
-       Compute the quantization loss $L_{\text{quant}}=\|z_{e}-z_{q}\|_{2}^{2}$;
-       Compute the reconstruction $\tilde{c}=f_{\text{dec}}(y,z_{q})$;
-       Compute the reconstruction loss $L_{\text{recon}}=\|c-\tilde{c}\|_{2}^{2}$;
-       GenCo Variant: Compute the optimization loss $L_{\text{opt}}=f_{\text{objective}}(g_{\text{solver}}(\tilde{c},y))$;
-       Compute the total loss: $L_{\text{total}}=L_{\text{recon}}+\beta_{1}L_{\text{quant}}$ + $\beta_{2}L_{\text{opt}}$;
-       Update the parameters of the encoder, decoder, and codebook to minimize $L_{\text{total}}$;
-end for
 Algorithm 4 Constrained Generator Training for VQVAE
+1:  Given: Training data distribution $D$ over problem info $y$ and known high-quality solutions $x$, regularization weight $\beta$, linear surrogate solver $g_{\text{solver}}$, nonlinear objective $f_{\text{objective}}$.
+2:  Output: Trained encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and codebook $E$
+3:  Initialize the parameters of the encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and the codebook $E=\{e_{1},e_{2},\ldots,e_{K}\}$ with $K$ embedding vectors
+4:  for $t=1$ to $T$ do
+5:     Sample $y,x$ from the distribution $D$
+6:     Compute the encoder output $z_{e}=f_{\text{enc}}(y,x)$
+7:     Find the nearest embedding vector $z_{q}=\arg\min_{e\in E}\|z_{e}-e\|_{2}^{2}$
+8:     Compute the quantization loss $L_{\text{quant}}=\|z_{e}-z_{q}\|_{2}^{2}$
+9:     Decode the embedding $\tilde{c}=f_{\text{dec}}(y,z_{q})$
+10:     Solve $\tilde{x}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$
+11:     Compute the reconstruction loss $L_{\text{recon}}=\|x-\tilde{x}\|_{2}^{2}$
+12:     Compute the optimization loss $L_{\text{opt}}=f_{\text{objective}}(\tilde{x})$
+13:     Compute the total loss: $L_{\text{total}}=L_{\text{recon}}+\beta_{1}L_{\text{quant}}+\beta_{2}L_{\text{opt}}$
+14:     Update the parameters of the encoder, decoder, and codebook to minimize $L_{\text{total}}$
+15:  end for